### v1

# On the Geometry of Semantics in Next-token Prediction

Abstract
Modern language models demonstrate a remarkable ability to capture linguistic meaning despite being trained solely through next-token prediction (NTP). We investigate how this conceptually simple training objective leads models to extract and encode latent semantic and grammatical concepts. Our analysis reveals that NTP optimization implicitly guides models to encode concepts via singular value decomposition (SVD) factors of a centered data-sparsity matrix that captures next-word co-occurrence patterns. While the model never explicitly constructs this matrix, learned word and context embeddings effectively factor it to capture linguistic structure. We find that the most important SVD factors are learned first during training, motivating using spectral clustering of embeddings to identify human-interpretable semantics, including both classical k-means and a new orthant-based method directly motivated by our interpretation of concepts. Overall, our work bridges distributional semantics, neural collapse geometry, and neural network training dynamics, providing insights into how NTP’s implicit biases shape the emergence of meaning representations in language models.
1 Introduction
Next-token prediction (NTP) is conceptually simple: predict the next word given a preceding sequence (aka context). Yet it trains models with remarkable ability to capture meaning.
A foundational step toward understanding how NTP leads to models acquiring semantic information is to first understand how explicit textual inputs—words and contexts—are represented in the models’ $d$-dimensional word and context vector representations (embeddings). Zhao et al. [2024] recently proved that in well-trained models, the geometry of these learned representations is characterized by the factorization of a data-sparsity matrix $\widetilde{\bm{S}}$ whose entries encode whether a particular word follows a given context in the training corpus.
While this characterization explains how models encode explicit training signals, it does not address the deeper question about semantic learning. Language conveys meaning through latent concepts that exist beyond explicit (context, word) pairs, leading us to investigate: How do NTP-trained models extract and encode these latent concepts? For instance, how do they capture semantic dichotomies like male/female, or grammatical categories like nouns and verbs? In other words, how does latent information shape the embedding geometry, and does this suggest ways to identify interpretable semantic structures in the embedding space? These questions are challenging because linguistic concepts are never explicitly given as inputs in the NTP objective. Yet, by analyzing the geometry of learned representations and its relation to the data-sparsity matrix, we show that latent linguistic structure emerges from the latter. Concretely, our contributions are summarized as follows.
Geometric emergence of latent concepts: We demonstrate that latent linguistic concepts learned by a large neural model through NTP emerge as principal components in the singular value decomposition of a centered data-sparsity matrix $\tilde{\mathbf{S}}$ (defn. in Sec. 3.1, see Fig. 1 for illustration). While $\widetilde{\bm{S}}$ is never explicitly formed during training, its structure naturally shapes the geometry of learned representations: concepts can be recovered as weighted combinations of word and context embeddings, where the weights are determined by the singular vectors of $\tilde{\mathbf{S}}$. This extends the characterization of the unconstrained-features model (UFM) by Zhao et al. [2024], revealing how embeddings inherently organize around latent concept dimensions.
Rate of learning: Singular values of the data-sparsity matrix quantify the significance of their corresponding concepts. By recognizing a connection between UFM with square-loss and the closed-form training dynamics of two-layer neural networks with orthogonal inputs derived by Saxe et al. [2013], we show that concepts associated with larger singular values are learned faster during training.
Semantics [^1] We broadly define “semantic” to encompass linguistic factors that humans consider when producing text, including grammatical (e.g., tense), syntactic (e.g., verb), and semantic (e.g., sentiment) elements. as orthant-based clustering: Not all concepts have human-interpretable meaning. Guided by the insight that top concepts are learned first during training, we find that human-interpretable semantic information is captured through specific positive/negative configurations (indicating how present/absent the specific concept is in that word or context) across a few top singular concept dimensions (see Fig. 2B for illustration). Identifying semantics through such sign configurations geometrically manifests as (spectral) clustering of word/context embeddings in orthant slices. Through experiments, we verify that this approach identifies distinct human-interpretable linguistic categories.
(A) Left Singular Matrix ($\bm{U}$)Refer to caption(B) Singular ValuesRefer to caption(C) Right Singular Matrix (${\bm{V}}$)Refer to caption
Figure 1: Illustration of semantic information in SVD of $\widetilde{\bm{S}}$ (rank=6). (A) Left singular matrix (U) where rows are analyzer vectors and columns are singular dimensions. The first dimension distinguishes between plants (“apple”, “tulip”, negative values) and animals (“dog”, “shark”, positive values). The second dimension further differentiates mammals (“dog”, “cat”, positive) from other animals (“trout”, “eagle”, negative). (B) Singular values in descending order of importance. (C) Right singular matrix (V) where the first dimension separates plant properties (“uses sunlight for food”) from animal properties (“hunts in groups”). The second dimension appears more complex: for animal properties, it contrasts mammal-specific traits (“stays warm-blooded”, positive) against aquatic traits (“lives in water”, negative), while for plant properties, it differentiates growth characteristics (“grows roots”, positive) from reproductive features (“reproduces with pollen”, negative).
Refer to caption (A) Embedding geometryRefer to caption (B) Orthant-based clusteringRefer to caption (C) Training data
Figure 2: NTP-trained word embeddings naturally organize into semantic clusters. (C) Training data (showing ${\bm{P}}^{\top}$) consists of pairs using the template “The organism that [attribute] is [subject]”, where attributes are rows and subjects are columns. (A) When trained with $d>V>rank(\widetilde{\bm{S}})=6$ and visualized in 3D PCA projection, both words (dots) and contexts (triangular marks) cluster by semantic category—animals (green/blue), and plants (red/orange). This demonstrates how semantic structure emerges naturally from the underlying data sparsity pattern, a process our work formalizes mathematically. (B) Visualization of orthant-based clustering and hierarchical semantic structure in a 3D subspace of context-analyzer vectors. Selected dimensions $\mathcal{K}=\{k_{1},k_{2},k_{3}\}={1,2,4}$ correspond to columns of ${\bm{V}}$ in Fig. 1C. Colored blocks represent different sign configurations: Green ($p=1$): $c_{k_{1}}=-1$ captures "animal" semantic. Blue ($p=2$): $[c_{k_{1}},c_{k_{2}}]=[-1,-1]$ refines to "mammal" semantic. Yellow ($p=3$): $[c_{k_{1}},c_{k_{2}},c_{k_{3}}]=[-1,-1,-1]$ further specifies "feline" semantic. This nested structure illustrates how increasing the number of active dimensions ($p$) yields progressively finer-grained semantic categories, where "mammal" is a subset of "animal", and "feline" is more specific within "mammal". Contexts shown with same color within each suborthant are example members according to Defn. 1.
Connecting distributional semantics and neural collapse (NC) geometries: On the one hand, linking to classical distributional semantics approaches based on heuristic forms of matrix factorization of data co-occurrences, we find using model abstractions popularized in the NC literature that NTP implicitly favors SVD factorization of the centered data-sparsity matrix. On the other hand, we broaden the scope of NC geometries of embeddings to encompass geometries of latent semantics – making the connection explicit, we show the emergence of semantics even for the default one-hot data setting in the NC literature.
2 Related Work
We bring together, in the context of NTP-trained language models, the following two classical ideas: (1) semantic patterns emerge from linear relationships between vector representations of textual inputs, and (2) these learned representations form geometries tied to language statistics via count matrices.
Semantics from Distributed Representations. The idea of extracting semantics from distributed representations, while rooted in earlier works Bengio et al. [2000], was popularized through word embedding research Mikolov et al. [2013a]. Two classically popular approaches include: (i) constructing semantic axes from centroids of pole words An et al. [2018], Fast et al. [2016], evolving from lexicon-based sentiment analysis Taboada et al. [2011], Hu and Liu [2004], and (ii) investigating semantic patterns through “word analogies,” which reveal that semantic relationships manifest as linear directions in embedding space Mikolov et al. [2013c], Levy et al. [2015], Drozd et al. [2016], Ethayarajh et al. [2019], Allen and Hospedales [2019]. In the context of modern LLMs, word analogies have been empirically observed Rezaee and Camacho-Collados [2022], Wijesiriwardene et al. [2024]. However, recent focus has shifted to techniques for distilling interpretable semantics through dictionary learning or sparse autoencoders Cunningham et al. [2023], Bricken et al. [2023], further showing these can be used for model steering Turner et al. [2023], Li et al. [2024] (techniques that in fact trace back to early work with static word embeddings, e.g., Murphy et al. [2012], Faruqui et al. [2015], Arora et al. [2018], Zhang et al. [2019]). Alternative approaches include geometric studies showing how linguistic features cluster in embedding spaces Coenen et al. [2019], Hewitt and Liang [2019]. Simultaneously, supervised methods such as linear probing have been employed to extract semantic and syntactic information from modern LLMs Alain and Bengio [2017], Hewitt and Manning [2019], Marks and Tegmark [2023], relying on the same principle of linear representation found in word analogies. Unlike these aforementioned works that empirically analyze trained models post-hoc, we focus on formalizing the mechanisms by which semantic structure emerges during NTP training. Our goal is not to introduce a state-of-the-art method for extracting semantics, but rather to understand, from an optimization viewpoint, how semantics emerge as a consequence of next-token prediction. While Park et al. [2023, 2024] have explored similar questions, here, rather than relying on a model for concept generation and its link to words/contexts, we directly analyze NTP.
Geometry from Co-occurrence Statistics. Vector space representations emerge from two approaches: (i) traditional count-based models using co-occurrence matrices, and (ii) prediction-based models that optimize token prediction objectives like Skip-gram or NTP. Levy and Goldberg [2014], Pennington et al. [2014] demonstrate that these approaches are fundamentally similar, as both probe underlying corpus co-occurrence statistics. Zhao et al. [2024] have extended this insight to causal NTP models with a key distinction: the co-occurrence matrix between contexts and words is inherently sparse, unlike the dense word-word co-occurrences in prior work. This distinction makes relevant recent advances in implicit regularization Ji and Telgarsky [2019], Ji et al. [2020] and neural collapse geometries Papyan et al. [2020b], Mixon et al. [2020], Fang et al. [2021], through which Zhao et al. [2024] analyze how word and context embeddings converge via an unconstrained features model, relating them to singular factors of the centered co-occurrence probability matrix’s support set (see Eq. (2)). We show that these singular factors encode rich semantic structure. While we argue that the natural language’s inherent multilabel nature is responsible for the emergence of rich semantics, making full circle to the neural-collapse literature, we show that the semantic interpretation applies to as simple a setting as one-hot imbalanced classification. Moreover, by connecting to the closed-form analysis of learning dynamics of linear neural networks by Saxe et al. [2013, 2019], Gidel et al. [2019], we show how singular values determine the learning order during training.
More details on related works deferred to App. E. See also Remark 1.
3 Background
Notations. For any integer $k$, $[k]:=\{1,\ldots,k\}$. Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$ respectively. For matrix $\bm{A}$, $\bm{A}[i,j]$ is its $(i,j)$-th entry, and for vector $\bm{a}$, $\bm{a}[i]$ is its $i$-th entry. $\mathds{1}$ is the all-ones vector and $\otimes$ is Kronecker product.
3.1 NTP Objective as Sparse Soft-Label Classification
[…]
$\mathcal{L}({\bm{W}},\bm{\theta})=\sum\nolimits_{j\in[m]}\hat{\pi}_{j}\cdot\ell\left({\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j});\hat{\bm{p}}_{j}\right)\,,$ (1)
where the loss $\ell$ penalizes deviations between the model’s logits ${\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j})$ for context $j$ and its corresponding soft-label vector $\hat{\bm{p}}_{j}$. Unless otherwise specified, we use the standard cross-entropy loss we set the loss $\ell$ to be the standard CE loss, i.e., $\operatorname{CE}\left(\mathbb{S}({\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j}))\,||\,\bm{p}_{j}\right)$, where $\mathbb{S}()$ denotes the softmax map of logits to the probability simplex.
For later use, define the support matrix $\bm{S}\in\{0,1\}^{V\times m}$ of the conditional probability matrix ${\bm{P}}=[\hat{\bm{p}}_{1},\ldots,\hat{\bm{p}}_{m}]\in\mathbb{R}^{V\times m}$. Formally, $\bm{S}[z,j]=1\,\Leftrightarrow\,{\bm{P}}[z,j]>0$. For context $j$, we refer to tokens $z$ for which $\bm{S}[z,j]=1$ as in-support tokens, and refer to all others as off-support. Central to our analysis is the centered support matrix
$\widetilde{\bm{S}}:=(\mathds{I}_{V}-\frac{1}{V}\mathds{1}_{V}\mathds{1}_{V}^{\top})\bm{S}\,.$ (2)
For convenience, we refer to $\widetilde{\bm{S}}$ as the data-sparsity matrix, though note that unlike the support matrix $\bm{S}$, its entries are not binary due to the centering operation:
$\widetilde{\bm{S}}[z,j]=\begin{cases}1-\nicefrac{{|{\mathcal{S}}_{j}|}}{{V}}&\text{if $\bm{S}[z,j]=1$}\,,\\ -\nicefrac{{|{\mathcal{S}}_{j}|}}{{V}}&\text{if $\bm{S}[z,j]=0$}\,.\end{cases}\,$
3.2 Geometry of Words and Contexts
Following Yang et al. [2017], we assume the model is sufficiently expressive to allow optimizing context embeddings in (1) freely, instead of abiding by their architecture-specific parameterization. Concretely, this modeling abstraction yields a simplified training objective
$\min\nolimits_{{\bm{W}},{\bm{H}}}~{}~{}\mathcal{L}({\bm{W}}{\bm{H}})+\frac{{\lambda}}{2}\|{\bm{W}}\|^{2}+\frac{{\lambda}}{2}\|{\bm{H}}\|^{2}\,,$ (NTP-UFM)
which jointly optimizes ${\bm{W}}\in\mathbb{R}^{V\times d}$ and ${\bm{H}}:=[\bm{h}_{1},\ldots,\bm{h}_{m}]\in\mathbb{R}^{d\times m}$, where $\bm{h}_{j}$ is the freely optimized embedding of context $\bar{\bm{x}}_{j}$. Since the minimization is unconstrained for both variables, we follow Mixon et al. [2022], Wojtowytsch [2021], Fang et al. [2021] in referring to this as the unconstrained features model (UFM) for NTP training. We refer to ${\bm{W}}$ and ${\bm{H}}$ as the word and context embedding matrices.
Eq. (NTP-UFM) includes ridge-regularization with weight $\lambda>0$. Zhao et al. [2024] have analyzed the geometry of solutions to (NTP-UFM) when $\lambda\rightarrow 0$. This limit (referred to in the literature as the regularization-path) serves as a proxy for the limiting behavior of gradient descent (GD) training as the number of iterations approaches infinity Ji et al. [2020]. For large embedding dimensions $d\geq V$,  [^2] See Sec. A.1 for a discussion on this assumption. as $\lambda\rightarrow 0$, Zhao et al. [2024] show the following:
1. Logits Convergence: The logit matrix $\bm{L}={\bm{W}}{\bm{H}}$ decomposes into two orthogonal components: one that grows unboundedly during training and another that remains finite. After normalization (required to characterize the convergence of diverging logits), only the unbounded component becomes dominant, denoted as $\bm{L}^{\rm{mm}}$. Crucially, $\bm{L}^{\rm{mm}}$ depends only on the data support matrix $\bm{S}$ (see Zhao et al. [2024] for details).
2. SVD factors of $\bm{L}^{\rm{mm}}$: Like logits, word and context embeddings grow unboundedly in magnitude, but converge in direction to ${\bm{W}}^{\rm{mm}}=\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ and ${\bm{H}}^{\rm{mm}}=\mathbf{R}^{\top}\bm{\Sigma}^{1/2}{\bm{V}}^{\top}$, respectively. Here, $\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ is the SVD of $\bm{L}^{\rm{mm}}$ and $\mathbf{R}$ is a partial orthogonal matrix.
3. Data-sparsity matrix as proxy: The matrix $\widetilde{\bm{S}}$ (see Eq. (2)) is a good proxy for $\bm{L}^{\rm{mm}}$.  [^3] This is formally proven in symmetric cases and supported by extensive empirical evidence. Put together, the singular factors of $\widetilde{\bm{S}}$ play a crucial role in determining the geometry of word and context embeddings learnt by NTP training. Denote this SVD as
$\widetilde{\bm{S}}:=\bm{U}\bm{\Sigma}{\bm{V}}^{\top},\quad$ (3)
where $\bm{U}\in\mathbb{R}^{V\times r}$, ${\bm{V}}\in\mathbb{R}^{m\times r}$ with $\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}\,,$ the singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered: $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}>0\,$, and $r\leq V-1$ denotes the rank.
4 Geometry of Concepts
4.1 Motivating Questions
While prior work explicitly describes the geometry of word and context embeddings, it leaves open two key questions: (i) What is the latent information that drives the word/context embeddings to this geometry? And, (ii) does the resulting geometry form structures (e.g. in the form of clusters) that lend themselves to human-interpretable semantics?
To illustrate this, Fig. 2A shows a 3D projection of trained word and context embeddings on a simplified dataset for visualization. The conditional probability matrix ${\bm{P}}$ is shown on the right. The geometry reveals clear structure: animal-related words and contexts cluster on the left, plant-related ones on the right. This suggests the emergence of latent concept information—specifically, a positive/negative semantic distinction.
From Sec. 3.2, we know word embeddings take the form ${\bm{W}}=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$, where $\bm{U}$ encodes the row-space of $\widetilde{\bm{S}}$ (Eq. (2)). Since rows of $\widetilde{\bm{S}}$ correspond to words, and words that appear in similar contexts (e.g., “the organism that grows roots is…” vs. “the organism that has green leaves is…”) have similar patterns, $\bm{U}$ may naturally encode semantic relationships. Its columns can be interpreted as a “concept basis”, where a word’s projection reflects its association with these latent concepts.
4.2 Principal Components as Concepts
We term each dimension $k\in[r]$ resulting from the SVD of $\widetilde{\bm{S}}$ as a concept. Adopting terminology from Saxe et al. [2019], we call the columns $\mathbf{u}_{k}\in\mathbb{R}^{V}$ and $\mathbf{v}_{k}\in\mathbb{R}^{m}$ of matrices $\mathbf{U}$ and $\mathbf{V}$ the “word analyzer vectors” and “context analyzer vectors,” respectively. These vectors represent the alignment of tokens and contexts with each concept: for word $z$ and concept $k$, the sign of $\mathbf{u}_{k}[z]$ indicates association ($\mathbf{u}_{k}[z]>0$) or opposition ($\mathbf{u}_{k}[z]<0$) to the concept, while its magnitude quantifies the strength of this relationship.  [^4] Note that each word analyzer vector $\bm{u}_{k},k\in[r]$ has both positive and negative entries (corresponding to association and opposition), because $\bm{u}_{k}$ is orthogonal to $\mathds{1}_{V}$ (which is in the nullspace of $\widetilde{\bm{S}}$). Words for which $|\bm{u}_{k}[z]|$ is small are neutral to this concept. The same interpretation applies to context components $\mathbf{v}_{k}[j]$, measuring each context’s alignment with the respective concept. Fig. 1 verifies the SVD of $\widetilde{\bm{S}}$ captures the dataset’s semantic information for the simplified [^5] This design isolates semantic concept extraction by minimizing grammatical influences. These simplifications are for visualization clarity only—the principles apply to complex datasets with combined grammatical/semantic attributes and to full autoregressive training (see Secs. 5 and A.2). dataset with syntax "The organism that [attribute] is [subject]" (see Fig. 2).
To represent concepts in the embedding space and thus directly characterize their relationships with word and context embeddings, we define the $d$-dimensional “word-concept representations” $\bm{u}^{\rm{d}}_{k}$ and “context-concept representations” ${\bm{v}}^{\rm{d}}_{k}$ for $k\in[r]$ as projections onto the respective spaces of words and contexts, that is (see also Sec. B for details), $\bm{u}^{\rm{d}}_{k}={\bm{W}}^{\top}\bm{u}_{k}$ and ${\bm{v}}^{\rm{d}}_{k}={\bm{H}}{\bm{v}}_{k}.$ This ensures that words or contexts more closely aligned with a specific concept have embeddings closer to the concept’s representation. Similar to classical semantic axes formed by averaging relevant words, these representations are weighted averages where weights ($\bm{u}_{k}$) reflect concept relevance. Fig. 10 visualizes two most significant concept representations and their relationships with word/context embeddings.
Remark 1.
The idea that SVD factors of a co-occurrence-type matrix convey latent semantic information dates back at least to latent semantic analysis of word-document co-occurrence matrices Landauer [1997], Deerwester et al. [1990] and later to pointwise mutual information matrices of word-word co-occurrences Levy and Goldberg [2014], Levy et al. [2015]. Beyond SVD factorization, researchers have explored alternative factorization techniques, such as non-negative matrix factorization Lee and Seung [1999], Ding et al. [2006] and sparse dictionary learning Murphy et al. [2012], Faruqui et al. [2015], and have applied them to various transformations of co-occurrence data, including modifications that address the sparsity of PMI matrices by eliminating negative entries (e.g., PPMI) or applying probability smoothing techniques Levy et al. [2015]. Our key conceptual contribution is that we let the training dynamics NTP optimization determine both the matrix representation of the data and the factorization method used to encode latent semantics: the model naturally utilizes the centered support matrix $\widetilde{\bm{S}}$ and its SVD factorization. We validate through structured examples in this section and through experiments in Sec. 5, that the model’s emergent representational choices successfully encode linguistic components without explicit engineering of the representation space.
4.3 Orthant-based Clustering in Singular Spaces
Concepts extracted from $\widetilde{\bm{S}}$, represent significant components from data. However, as noted, for example, by Chersoni et al. [2021], Piantadosi et al. [2024], not all concepts individually correspond to linguistically interpretable factors. Instead, we posit that human-interpretable semantics can be identified by specific combinations of several concepts. Here, we formalize this hypothesis by relating it to an orthant-based clustering of word/context representations.
Recall from Sec. 4.2 that the $k$-th word analyzer vector $\bm{u}_{k}$ categorizes words $z$ based on $\texttt{sign}(\bm{u}_{k}[z])$: words with large $|u_{k}[z]|$ exhibit strong correlation with concept $k$, with $\texttt{sign}(\bm{u}_{k}[z])$ indicating whether this correlation is positive/negative. We extend this categorization beyond single dimensions to multiple concept dimensions simultaneously.
Let $\mathcal{K}=\{k_{1},\ldots,k_{p}\}$ denote a set of $p\leq r$ selected dimensions $k_{i}\in[r],i\in[p]$. Given $\mathcal{K}$, we define $2^{p}$ possible signature configurations $C=C(\mathcal{K})=[c_{k_{1}},\ldots,c_{k_{p}}]\in\{\pm 1\}^{p}$, where $c_{k_{i}}\in\{\pm 1\}$ indicates whether dimension $k_{i}$ contributes positively or negatively to the semantic category. Each signature configuration represents a potential semantic category, characterized by its member words and their degrees of association with the category.
Definition 1.
For $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ with sign configurations $c_{k_{i}}\in\{\pm 1\}$ for dimensions $k_{i}\in[r]$:
  • A word $z$ is a member of $C$ (denoted $z\in C$) if and only if $\texttt{sign}(\bm{u}_{k_{i}}[z])=c_{k_{i}}$ for all $k_{i},i\in[p]$.
  • The typicality of a member $z\in C$ is defined as: $\text{Typicality}(z;C)=\sum\nolimits_{i\in[p]}\big{|}\bm{u}_{k_{i}}[z]\big{|}\,.$
Analogous definitions hold for contexts.
Geometrically, each configuration corresponds to a $p$-dimensional orthant (or orthant sheet) in the $d$-dimensional embedding space. Membership indicates which word/context embeddings lie within this orthant, effectively creating an orthant-based clustering in the subspace spanned by the selected concept dimensions $\mathcal{K}$. The typicality of a member measures the L1-norm (other choices are also possible) of its representation’s projection (the $z$-th row of $\bm{U}$ for words, $j$-th row of ${\bm{V}}$ for contexts) onto the selected concept dimensions. Fig. 2B and Sec. 5 demonstrate that this orthant-based clustering reveals interpretable semantic categories, while also inducing a hierarchical structure, where the number of concept dimensions determines the semantic granularity: broad linguistic categories (e.g., verbs) require fewer dimensions, while more specific categories (e.g., past-tense verbs) need more dimensions for precise characterization. See also Sec. C and Fig. 11 in the appendix.
4.4 Rate of Learning
We show that the hierarchical structure of semantic concepts discussed in Sec. 4.3 is reflected in their emergence during training: broad semantic distinctions are learned before finer-grained features. Specifically, we track the singular values of the logit matrix $\bm{L}_{t}={\bm{W}}_{t}{\bm{H}}_{t}$ at training iteration $t$ when minimizing the unconstrained (${\lambda}\rightarrow 0$) objective (NTP-UFM). On the one hand, at the end of training ($t\rightarrow\infty$), singular values exhibiting the largest divergence correspond to the largest singular values of $\widetilde{\bm{S}}$ Zhao et al. [2024], which represent the most important concepts. On the other hand, at initialization, all singular values of $\bm{L}_{t}$ are small (following typical small-scale initialization). As training progresses, Fig. 18 shows that the singular values increase at different rates: those that ultimately achieve the greatest magnitudes begin diverging earlier.
Fig. 3 provides empirical evidence of this hierarchical concept acquisition. The confusion matrices from a 2-layer transformer trained on natural language data reveal that the model first masters broad categorical distinctions (e.g., "Animals" vs. "Plants") before progressing to more specific semantic features. This progression hints at a coarse-to-fine learning order, which we formalize in the following theoretical analysis.
True ClassEpoch 0Refer to captionEpoch 20Refer to captionEpoch 40Refer to captionEpoch 200Refer to captionPredicted Class
Figure 3: Confusion matrices showing the progression of concept learning in a 2-layer transformer (embedding dimension $d=128$) across training epochs (0, 20, 40, 200). Each matrix cell shows the number of predictions for each true class. Classes correspond to support sets defined in Fig. 2C, with semantic labels (e.g., "Animals", "Plants") given in brackets where applicable. For example, the support set dog, wolf, cat, tiger, sparrow, eagle, trout, shark represents the "Animals" class (first row/column). The matrices illustrate a temporal learning hierarchy: broad categories like "Animals" and "Plants" emerge by epoch 20, while finer-grained concepts (e.g., "Reproduces with pollen", "Lives in water") develop later, between epochs 40–200. This reflects that high-variance (large singular value) concepts—often aligned with broader semantics—are learned earlier during training.
To gain theoretical insights into this behavior, we analyze (NTP-UFM) with square loss. [^6] This choice is motivated by the technical challenges in finite-time characterization of GD dynamics for CE loss due to coupling between singular directions. The simplified, square loss, could be seen as a first-order approximation of CE loss around zero initialization or in the high-temperature limit. Also, while CE loss is predominant in practice, we note that square-loss-type objectives have demonstrated practical success, as in the GloVe model Pennington et al. [2014], which achieves performance comparable to word2Vec Mikolov et al. [2013d]. We hope that future work can prove the analogous phenomena we observe empirically for CE loss. Specifically, we show in Sec. D in the appendix that the NTP-UFM setting is equivalent to the setting studied extensively by Saxe et al. [2013, 2019], Gidel et al. [2019]; thus, with sufficiently small initialization and learning rate, the $k$-th singular value of $\bm{L}_{t}$ undergoes a sharp phase transition from (almost) 00 to its limit $\sigma_{k}$ at time $T_{k}\propto 1/\sigma_{k}$ (see Thm. 1). In other words, more significant concepts (corresponding to larger singular values) are learned first, i.e. $T_{1}\leq T_{2}\leq\cdots\leq T_{r}$.
We further validate this behavior in a controlled imbalanced setting with one-hot labels. Two majority classes each appear across many contexts, while two minority classes appear in only one. As shown in Figure 9, the model initially predicts all inputs as the dominant majority class (Step 0), then begins separating the majority classes (Step 13), followed by gradual recognition of the minority classes (Step 30), and resolves confusion between them later in training (Step 55). This progression reflects the emergence of concepts in order of their singular values, consistent with theoretical analysis.
5 Experiments
5.1 Setup
Datasets. We conduct experiments on both synthetic data and pretrained models. For our synthetic experiments, we constructed two datasets by subsampling from two distinct corpora: TinyStories Eldan and Li [2023] and WikiText-2 Merity et al. [2017]. For both datasets, we fix the vocabulary size to $V=1,000$ and extract $m=10,000$ contexts, creating the "Simplified TinyStories/WikiText" datasets. We extracted the contexts by sampling frequent contexts of lengths 2-6 from the entire datasets. This variation in context-length emulates the autoregressive nature of training with sliding window, which yields rich data-sparsity patterns that in turn result in richer semantic information (compared to the last-token visuals in Sec. 4). For our experiments on pretrained transformers, we chose GPT-2 Radford et al. [2019] and extracted $50,000$ text sequences, each 11 tokens in length from GPT-2’s training corpus, the OpenWebText dataset. By segmenting these sequences into contexts of varying lengths (1-10 tokens), we generated $500,000$ contexts, which we designate as the "simplified OpenWebText" dataset.
Concepts and semantics. Following Sec. 4, for the simplified datasets where it is computationally feasible, we construct the $V\times m$ data-sparsity matrix $\widetilde{\bm{S}}$ and compute its SVD. The left/right singular vectors serve as our word/context concepts. As per Sec. 4.3, we also consider signature-based concept combinations, which we examine for semantic interpretability as discussed in Visuals below. For completeness, we also experiment with $k$-means spectral clustering as an alternative way to derive interpretable semantics from top important concepts. For the OpenWebText dataset, instead of performing SVD on $\widetilde{\bm{S}}$ (which is computationally expensive to construct and factorize), we work directly with the embeddings of the pretrained model. For word concepts, we extract ${\bm{W}}$ from GPT-2’s decoder and take the matrix $\bm{U}$ of word concepts to be its left singular matrix. For context concepts, we form a $d\times m$ matrix ${\bm{H}}$ by concatenating the embeddings (last-MLP layer representations before the decoder) of GPT-2 for each context in our subset. This matrix corresponds to only a portion of the “true” embedding matrix for the entire dataset. To ensure the coupling between ${\bm{W}}$ and ${\bm{H}}$ that is inherent to the SVD of $\widetilde{\bm{S}}$ (which we don’t have direct access to), we compute the matrix ${\bm{V}}$ of context concepts as ${\bm{V}}={\bm{H}}^{\top}\mathbf{R}\bm{\Sigma}^{-1/2}$, where $\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ denotes the SVD of ${\bm{W}}$ that we already computed.
Visuals. To interpret the semantic information, we visualize the top $40$ words/contexts associated with each concept. “Top” here is measured with respect to the typicality score in Defn. 1. For visualization, we use word clouds where size is proportional to its typicality score. We often use the pound sign (#) to indicate token start.
5.2 Findings
Individual concepts are not interpretable. We have seen in Sec. 4 that not all concepts individually correspond to linguistically interpretable factors. For completeness, we verify this conclusion for the Simplified TinyStories/WikiText datasets in Fig. 14 in the appendix. This observation is well-reported in the literature, e.g., Chersoni et al. [2021], and is also related to the recently popularized superposition hypothesis Elhage et al. [2022].
Combining concepts with signature-configurations reveals semantics. We demonstrate that combining non-interpretable concepts can reveal human-interpretable semantic information. Here, we show this by implementing the idea of orthant-based clustering of Sec. 4.3. Specifically, we combine $p=5$ to $7$ concepts in successive order starting from the most important one, i.e. $\mathcal{K}=\{1,\ldots,p\}$. Fig. 4 shows human-interpretable semantics identified by orthant-based clustering in Simplified TinyStories and WikiText-2 datasets. In Fig. 11, we show how orthant-based clustering progressively reveals hierarchical semantic structure as we combine more concept dimensions, illustrated through the example of verb forms organized by grammatical distinctions. Starting with a single dimension ($\mathcal{K}=\{1\}$), the categories remain broad and difficult to interpret. However, as we incorporate additional dimensions, increasingly refined semantic categories emerge, ultimately distinguishing between different verb forms such as past/present tense, and present continuous forms.
Past-tense verb
Present-tense verb
Preposition
Proper names
Month
Number
Unit
Preposition/Conjunction
Figure 4: Semantic clusters based on signature configurations of Sec. 4.3. Top: Simplified TinyStories dataset showing Past-tense verbs, Present-tense verbs, Prepositions, and Proper names. Bottom: Simplified Wikitext dataset showing Months, Numbers, Units, and Prepositions/Conjunctions. Each configuration is shown in its title.
Past-tense verbs Refer to caption
Proper names Refer to caption
Numbers Refer to caption
Conjunctions Refer to caption
Figure 5: $k$-means cluster with top-$p=\log_{2}(k)$ dimensions of word-analyzer vectors. Left: 2 of 32-means clusters with $\mathcal{K}=\{1,\ldots,5\}$ on Simplified TinyStories representing Past-tense verbs and Proper names; Right: 2 of 64-means cluster with $\mathcal{K}=\{1,\ldots,6\}$ on simplified WikiText, representing Numbers and Conjunctions.
$k$-means spectral clustering. As a complementary approach to our orthant-based clustering, we also investigate k-means spectral clustering. Specifically, we perform k-means clustering according to pairwise distances of analyzer vectors with their top $p=\log_{2}(k)$ dimensions (corresponding to top-$p$ most important concepts). Vanilla $k$-means spectral clustering selects $k$ principal dimensions. The modification here is inspired by the orthant-based clustering for which $p$ concepts correspond to $2^{p}$ orthants. Fig. 5 shows semantic meaning in a subset of the recovered clusters with $k=32$ and $k=64$ for Simplified TinyStories and WikiText-2, respectively; see Sec. F.5 for full results. See also discussion in Sec. A.3.
Relation of words and context semantics. Fig. 2A demonstrates a relationship between context embeddings and word embeddings that share semantic features: contexts and words representing the same semantic concept (e.g., plant or animal) cluster proximally in the embedding space. This pattern extends to the richer datasets we study here. To verify this, we perform orthant-based clustering of contexts and words based on the same signature configuration of a subset of concept dimensions. Fig. 6 shows this for Simplified TinyStories: contexts containing modal constructions like "want/like/love/have to" correspond strongly with word clusters containing infinitive verbs. Similarly, contexts with intensifier patterns "was so/very" align with clusters of descriptive adjectives. This relation reflects a natural dual semantic interpretation of contexts: they can be understood either through the direct semantic meaning of their constituent words or through the semantic properties of words that typically follow them. The latter interpretation aligns naturally with the predictive nature of NTP, where contexts acquire meaning partly from what they predict.
  Refer to caption Refer to caption
Figure 6: Context-word cluster pairs sharing identical sign patterns across top 5 concepts. Left and right pair illustrates the “verb+to infinitive” and “was so/very + adj” structures, confirming semantic alignment between contexts and their in-support words.
Semantics in pretrained models. We find interpretable semantics from simplified datasets are primarily grammatical or syntactic due to limited sample size and context simplicity. While these features make analytical computations feasible (e.g., computing/factorizing $\widetilde{\bm{S}}$), they compromise the richness of sparsity patterns and, in turn, the emergence of semantics. To verify that richer sparsity patterns yield richer semantics, we additionally experiment with pre-trained GPT-2. See Fig. 7 and Sec. F.2 in appendix for additional results.
Positive
Negative
Chemical
Medical
Coorperation
Digital entertainment
Cultural
Societal
Figure 7: Semantics identified by orthant-based clustering on GPT2’s word embeddings.
6 Conclusion and Limitations
While continuing to advance LLMs’ capabilities, understanding their inner workings is crucial. We study how NTP optimization guides the geometry of word and context representations to organize around latent semantic information. Without explicitly computing it, the model encodes such information in the SVD factors of an implicit data-sparsity matrix. As an early contribution in this direction, our analysis motivates several future directions: (1) While the assumption $d\geq V$ allows for rich geometric structures, a formal investigation is needed for the $d<V$ case prevalent in practice (see Sec. A.1). (2) Our analysis assumes sufficient expressivity and training time, which may become increasingly relevant with improved compute or in data-limited regimes, but relaxing them or experimentally validating their applicability in the spirit of our GPT-2 experiments (e.g., in the spirit of Wu and Papyan [2024]) is important. (3) Understanding how transformer architectures specifically fit into our analytical framework remains an open challenge. (4) Some conclusions rely on square loss and our work highlights a significant gap in understanding CE loss dynamics, even in simple bilinear models. (5) While our approach differs from concurrent efforts to explain semantic emergence in LLMs via probing or sparse-dictionary learning, future work could explore connections and potential applications to model improvement techniques.
7 Acknowledgments
This work was funded by the NSERC Discovery Grant No. 2021-03677, the Alliance Grant ALLRP 581098-22, and an Alliance Mission Grant. The authors also acknowledge use of the Sockeye cluster by UBC Advanced Research Computing.
Appendix A Additional Discussions
A.1 On the Assumption $d\geq V$
Our analysis has assumed the embedding dimension $d$ is at least as large as the rank of the data-sparsity matrix (guaranteed when $d\geq V$), allowing word and context embeddings to form geometric structures representing all concepts encoded in the matrix’s singular factors Zhao et al. [2024]. As noted in Zhao et al. [2024], while this assumption differs from current practice in LLMs, the geometry of embeddings remains rich in this setting. Importantly, this assumption is less restrictive than requiring $d>C$ in one-hot classification with $C$ classes. There, due to collapse of embeddings from the same class, $d>C$ effectively requires the dimension to exceed the number of training examples. In contrast, for NTP, the number of contexts $m$ can be (and typically is) much larger than $d$, allowing for rich geometric arrangements of context embeddings in the lower-dimensional space and non-trivial emergence of concepts, which is confirmed throughout our case studies.
That said, the current trend of modern language models typically employing $d<V$, raises the question: Which concepts are prioritized when the model cannot capture all of them? We hypothesize that during NTP training, models learn to represent the $d$ most significant concepts, corresponding to the largest singular values of the data-sparsity matrix. While this requires a separate study that is beyond our current scope, our preliminary investigation using NTP-UFM with square-loss on a synthetic dataset supports this hypothesis. Fig. 19 shows the learned logit matrix’s singular values converging to the $d$ largest singular values of the data-sparsity matrix during training. Following Sec. 4.4, we used square loss rather than CE to ensure bounded singular values. While a rigorous analysis for $d<V$ remains future work, since meaningful semantic categories emerge from combinations of concepts, even a reduced set of core concepts could enable rich semantic representations through the orthant clustering mechanism of Sec. 4.3.
A.2 Role of Autoregression
We have shown that concepts emerge during NTP training as principal directions of $\widetilde{\bm{S}}$, a centered version of the support matrix $\bm{S}$. Each column $j$ of $\bm{S}$ corresponds to a context and can be viewed as a binary (multi-)label vector $\bm{s}_{j}$, where entries of $1$ indicate tokens that appear as next-tokens for that context in the training data. While concepts are determined by the principal directions of this label matrix, this might seem limiting: contexts can relate not only through shared next-tokens but also through their intrinsic structure (e.g., overlapping constituent tokens).
The autoregressive nature of training naturally captures these structural relationships as follows. As the model processes labels for progressively longer contexts, each context $(z_{1},\ldots,z_{t})$ contributes to concept formation both through its distribution over next-tokens $z_{t+1}$ and through the labels of its shorter subsequences $(z_{1},\ldots,z_{t-1})$, $(z_{1},\ldots,z_{t-2})$, etc.—each representing distinct columns in $\bm{S}$. This overlapping window structure produces fine-grained label information, creating rich sparsity patterns in $\bm{S}$ (and consequently in $\widetilde{\bm{S}}$) that yield nontrivial concepts with varying significance, as reflected in word and context embeddings.
A.3 Relationship to Spectral Clustering
The orthant-based clustering approach introduced in this work shares foundational similarities with spectral clustering while providing distinct advantages for linguistic interpretation. Both methods operate in reduced-dimensional spaces and leverage embedding geometry to identify meaningful clusters. First, while both approaches leverage singular factors, the orthant-based method provides a more direct link (in view of concept definition in Sec. 4) between NTP and the emergence of semantic structure by explicitly interpreting the sign patterns of SVD dimensions as semantic indicators.
An interesting property of our method is its selective nature: while $p$ dimensions generate $2^{p}$ potential orthants, only a subset contains semantically coherent word groupings. This selectivity reflects linguistic reality, where not all mathematical combinations yield meaningful semantic categories. In contrast, spectral clustering typically partitions the space completely, which may not always align with natural semantic boundaries in language. Finally, the orthant-based method offers semantic interpretation of membership and typicality, and dimensionality control through parameter $p$. This aligns with language model training dynamics, as the most important singular value components are learned first during training, making the resulting clusters reflective of the model’s knowledge acquisition process. To this respect, note that unlike orthant-based clustering producing a potential maximum of $2^{p}$ clusters when selecting $p$ top concepts, typical clustering algorithms such as vanilla $k$-means on principal components would result in $p$-clusters with the same number of selected concepts. Thus, when we evaluate spectral methods in our experiments, we choose $k$ on the order of $2^{p}$ for the value of $p$ that we find orthant-based clustering reveals semantically informative orthants.
A.4 Neural Collapse Meets Semantics
The emergence of concepts with rich semantic meanings stems from the interplay of labels (encoded in $\bm{S}$ or its centered version $\widetilde{\bm{S}}$) across different contexts. Consider the minimal label richness case: each context has exactly one next-token, with contexts distributed equally across the vocabulary. Here, $\bm{S}$ can be rearranged as $\mathds{I}_{V}\otimes\mathds{1}_{m/V}^{\top}$, yielding trivial concepts where all singular directions contribute equally. This setting parallels standard balanced one-hot classification studied in neural collapse literature, where last-layer embeddings and weights form highly symmetric aligned structures Papyan et al. [2020a], reflecting the symmetric nature of $\bm{S}$ where labels induce no interesting conceptual structure.
However, such balanced settings never occur in autoregressive NTP, which as discussed in Sec. A.2, produces rich label formations. In these richer settings, $\bm{S}$ induces a geometry of concepts that provides semantic interpretation to the embedding geometry studied in the neural collapse literature. To illustrate this new perspective, we demonstrate that meaningful concepts emerge even in one-hot classification—the traditional focus of neural collapse literature—given minimal deviation from perfect balance. Specifically, consider STEP imbalances with ratio $R$: $V/2$ majority classes each have $R>1$ times more samples than minority classes.
Here, Thrampoulidis et al. [2022] shows the learned logit matrix converges to $\widetilde{\bm{S}}$, whose singular values exhibit a three-tier structure:
$\sigma_{1}=\ldots=\sigma_{V/2-1}=\sqrt{R}>\sigma_{V/2}=\sqrt{(R+1)/2}>\sigma_{V/2+1}=\ldots=\sigma_{V}=1.$ The left singular vectors matrix $\bm{U}$ takes a sparse block form:
$\bm{U}=\left[\begin{array}[]{ccc}\mathbb{F}&-\sqrt{\frac{1}{V}}\mathbf{1}&\mathbf{0}\\ \mathbf{0}&\sqrt{\frac{1}{V}}\mathbf{1}&\mathbb{F}\end{array}\right]\in\mathbb{R}^{V\times(V-1)}\,,$
where $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ is an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$. [^7] For concreteness, $\mathbb{F}$ can be constructed using the discrete cosine transform matrix, excluding the constant column: $\mathbb{F}[i,j]=\sqrt{\frac{4}{V}}\cdot\cos\left(\frac{\pi(2i-1)j}{V}\right)$ for $i\in[V/2]$, $j\in[V/2-1]$ The structure of $\bm{U}$ reveals three distinct types of concepts, corresponding to the three tiers of singular values: (1) First $V/2-1$ columns (non-zero only for majority classes) represent distinctions among majority classes; (2) Middle column (singular value $\sqrt{(R+1)/2}$) has opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy; (3) Last $V/2-1$ columns (non-zero only for minority classes) capture distinctions among minority classes. This hierarchical structure shows the network learns concepts in order: first majority class distinctions, then the majority-minority split, and finally minority class differences. See Fig. 13 for a visualization of $\bm{U}$ and Fig. 9 for an experiment that confirms the above semantic interpretation of concepts (columns of $\bm{U}$).
True ClassStep 0Refer to captionmaj 1maj 2min 1min 2maj 1maj 2min 1min 2Step 13Refer to captionmaj 1maj 2min 1min 2Step 30Refer to captionmaj 1maj 2min 1min 2Step 55Refer to captionmaj 1maj 2min 1min 2Step 60 Refer to captionmaj 1maj 2min 1min 2Predicted Class
Figure 9: Experimental illustration of the fact that important concepts are learned first. Confusion matrix evolution during training of a 3-layer convolutional network on an imbalanced MNIST data. The matrices show five snapshots at different training steps (0, 13, 30, 55, 65) for four classes: two majority classes (Maj1, Maj2) with 100 samples each and two minority classes (Min1, Min2) with 10 samples each. Training progresses from left to right: initially classifying all data as Maj1 (Step 0), then correctly identifying majority classes while misclassifying minority classes (Step 13), gradually improving minority class recognition (Step 30), confusion between minority classes only (Step 55), and finally achieving perfect classification by Step 60.
Appendix B From Sparsity Language Pattern to Concepts Geometry
Recall the centered data-sparsity matrix $\widetilde{\bm{S}}$ and its SVD
$\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top},\quad\text{where}\quad\bm{U}\in\mathbb{R}^{V\times r},\bm{\Sigma}\in\mathbb{R}^{r\times r},{\bm{V}}\in\mathbb{R}^{m\times r}\quad\text{and}\quad\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}\,,$
and the singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered:
$\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}>0\,.$
Adopting terminology from Saxe et al. [2019], we interpret the columns $\bm{u}_{k}\in\mathbb{R}^{V},{\bm{v}}_{k}\in\mathbb{R}^{m},k\in[r]$ of $\bm{U},{\bm{V}}$ as word and context analyzer vectors for concept $k$. For each word $z\in\mathcal{V}$ and each word-concept $k\in[r]$, the component $\bm{u}_{k}[z]$ represents how present or absent is a word $z$ in context $k$. Respectively for contexts.
Specifically, we think of column dimensions of $\bm{U}$ as semantic dimensions that capture semantic categories.
Q: How do we define word-concept and context-concept representations, i.e. $d$-dimensional representations of word and context analyzer vectors for various concepts?
Let ${\bm{W}}\in\mathbb{R}^{V\times d}$ and ${\bm{H}}\in\mathbb{R}^{d\times m}$ be the representations of words and contexts. We then define word-concept representations $\bm{u}^{\rm{d}}_{k}$ and context-concept representations ${\bm{v}}^{\rm{d}}_{k}$ for $k\in[r]$ as projections onto the spaces of word and context representations, respectively. Specifically, for projection matrices
$\mathds{P}_{\bm{W}}={\bm{W}}^{\top}({\bm{W}}{\bm{W}}^{\top})^{-\dagger}{\bm{W}}\qquad\text{and}\qquad\mathds{P}_{\bm{H}}={\bm{H}}({\bm{H}}^{\top}{\bm{H}})^{-\dagger}{\bm{H}}^{\top}\,,$
let
$\bm{u}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{W}}{\bm{W}}^{\top}\bm{u}_{k}\,$
${\bm{v}}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{H}}{\bm{H}}{\bm{v}}_{k}\,.$
Let’s now simplify these by using the known SVD representation of ${\bm{W}}$ and ${\bm{H}}$. Using this representation (i.e. use ${\bm{W}}\leftarrow{\bm{W}}^{\rm{mm}}$, ${\bm{H}}\leftarrow{\bm{H}}^{\rm{mm}}$) we compute
$\mathds{P}_{\bm{W}}=\mathbf{R}\mathbf{R}^{\top}=\mathds{P}_{\bm{H}}.$
Thus,
$\bm{u}^{\rm{d}}_{k}$ $=\mathbf{R}\mathbf{R}^{\top}\mathbf{R}\sqrt{\bm{\Sigma}}\bm{U}^{\top}\bm{u}_{k}=\sigma_{k}\mathbf{R}\bm{e}_{k}={\bm{W}}^{\top}\bm{u}_{k}\,$ (4a)
${\bm{v}}^{\rm{d}}_{k}$ $=\sigma_{k}\mathbf{R}\bm{e}_{k}={\bm{H}}{\bm{v}}_{k}=\sum_{j\in[m]}{\bm{v}}_{k}[j]\cdot\bm{h}_{j}\,.$ (4b)
We conclude that the $d$-dimensional representations of word and context analyzer vectors are the same. This is intuitive since concepts are categories in a certain sense broader than words/contexts, where the latter can be thought of as realizations of the concept in the form of explicit constituents of natural language. We thus refer to $\bm{u}^{\rm{d}}_{k}={\bm{v}}^{\rm{d}}_{k}=\mathbf{R}\bm{e}_{k}$ as the representation of concept $k$. See Fig. 10 for a visualization.
Finally, observe from Eq. (4) that the $k$-th concept representation is given by a weighted average of word or context embeddings with weights taken by the resppective context analyzer vectors.
Figure 10: Left: Context embeddings; Right: Word embeddings. In both plots, concept representations (blue, green) are computed using Eq. (4). Projections onto these concept axes (${\bm{v}}^{\rm{d}}_{1}$, ${\bm{v}}^{\rm{d}}_{2}$) quantify semantic relevance. The first concept (${\bm{v}}^{\rm{d}}_{1}$, blue line) defines a spectrum from plant-related to animal-related elements, with animals (“hunts in groups”, “dog”) projecting positively and plants (“uses sunlight for food”, “apple”) projecting negatively along the same axis. The second concept (${\bm{v}}^{\rm{d}}_{2}$, green line) represents a continuum from non-mammalian to mammalian traits, with mammalian words and features (“cat”, “stays warm-blooded”) having positive projections and non-mammalian ones (“trout”, “has feathers”) having negative projections. Please note that the concepts are not orthogonal in the plots because it is PCA projection to 2D space.
Appendix C On the hierarchical structure of language
As discussed in Sec. 4.3, language semantics exhibit a hierarchical structure, with more comprehensive semantics involving numerous concepts being inherently more detailed. Each concept’s significance is indicated by its associated singular values; higher values denote concepts critical for differentiating words or contexts. Moreover, as discussed in Sec. 4.4, broader semantic categories are acquired more rapidly, indicating a sequential learning progression from general to more specific concepts. Fig. 11 below provides an illustration of this hierarchical structure.
Figure 11: Hierarchical semantic structure revealed by orthant-Based clustering on Simplified TinyStories. Top: Hierarchical semantic structure emerges as we combine concepts. Each row represents clusters formed by combining an increasing number of concept dimensions (from 1 to 5), with each word cloud displaying the most typical words for a specific sign configuration. Semantic categories become progressively more fine-grained moving downward as additional concept dimensions refine the classifications. Bottom: Detailed view of verb semantics. The combination of the first two concept dimensions ($\mathcal{K}=\{1,2\}$ with configuration $C_{\mathcal{K}}=\{+1,+1\}$) identifies verbs. Adding the third dimension further distinguishes between past tense ($C_{\mathcal{K}}=\{+1,+1,+1\}$) and present tense ($C_{\mathcal{K}}=\{+1,+1,-1\}$). The present tense category is further refined by the fifth dimension to differentiate present action verbs from present continuous (-ing) forms. This hierarchical organization mirrors linguistic grammar structures and supports our hypothesis that language models learn fundamental grammatical distinctions (verb vs. non-verb) before more specific ones (tense and aspect distinctions).
Appendix D Connections to closed-form training dynamics Saxe et al. [2013]
Consider the following square-loss NTP-UFM proxy:
$\min_{{\bm{W}},{\bm{H}}}~{}~{}\left\{\sum_{j=1}^{m}\|\widetilde{\bm{s}}_{j}-{\bm{W}}\bm{h}_{j}\|^{2}=\|\widetilde{\bm{S}}-{\bm{W}}{\bm{H}}\|^{2}\right\}\,.$ (5)
which fits logits ${\bm{W}}{\bm{H}}$ to the centered sparsity matrix $\widetilde{\bm{S}}$. For one-hot labels, this reduces to standard square-loss classification, which has shown competitive performance to CE minimization in various settings (Hui and Belkin [2020], Demirkaya et al. [2020]). However, in our soft-label setting, the choice of loss function requires more careful consideration. While one could follow Glove’s approach Pennington et al. [2014] by using $\log({\bm{P}})$ instead of $\widetilde{\bm{s}}$, this creates issues with ${\bm{P}}$’s zero entries. Since CE loss minimization leads ${\bm{W}}$ and ${\bm{H}}$ to factor $\widetilde{\bm{s}}$, we thus maintain $\widetilde{\bm{s}}$ as the target in (5).
The neural-collapse literature has extensively studied Eq. (5) for one-hot $\widetilde{\bm{s}}$, primarily in the balanced case (e.g., Mixon et al. [2020], Han et al. [2021], Súkeník et al. [2023], Tirer and Bruna [2022]) but recently also for imbalanced data (e.g., Liu [2024], Hong and Ling [2023]). Most works focus on global minima of regularized UFM, with less attention to unregularized cases or training dynamics. While some landscape analyses provide partial answers about global convergence (e.g., ), they are limited to regularized cases and don’t characterize dynamics. Even for square loss, where Mixon et al. [2022], Han et al. [2021] study training dynamics—notably Han et al. [2021]’s analysis of the ’central path’ in balanced one-hot cases—these results rely on approximations. Thus, a significant gap remains in understanding UFM training dynamics, even for simple balanced one-hot data with square loss.
By interpreting the UFM with square loss in Eq. (5) as a two-layer linear network with orthogonal inputs, we identify an unexplored connection to Saxe et al. [2013], Gidel et al. [2019]’s analysis. They provide explicit characterization of gradient descent dynamics (with small initialization) for square-loss UFM. The key insight is rewriting (5) as $\sum_{j\in[m]}|\widetilde{\bm{s}}_{j}-{\bm{W}}{\bm{H}}\bm{e}_{j}|^{2}$ with orthogonal inputs $\bm{e}_{j}\in\mathbb{R}^{m}$. This enables direct application of their result, originally stated in Saxe et al. [2013] and formalized in Gidel et al. [2019]. For completeness, we state this here in our setting and terminology as a corollary below.
Proposition 1.
Consider gradient flow (GF) dynamics for minimizing the square-loss NTP-UFM (5). Recall the SVD $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$. Assume weight initialization
${\bm{W}}(0)=e^{-\delta}\bm{U}\mathbf{R}^{\top}\quad\text{and}\quad{\bm{H}}(0)=e^{-\delta}\mathbf{R}{\bm{V}}^{\top}$
for some partial orthogonal matrix $\mathbf{R}\in\mathbb{R}^{d\times r}$ ($\mathbf{R}^{\top}\mathbf{R}=\mathds{I}_{r}$) and initialization scale $e^{-\delta}$. Then the iterates ${\bm{W}}(t),{\bm{H}}(t)$ of GF are as follows:
${\bm{W}}(t)=\bm{U}\sqrt{\bm{\Sigma}}\sqrt{{\bm{A}}(t)}\mathbf{R}^{\top}\qquad\text{and}\qquad{\bm{H}}(t)=\mathbf{R}\sqrt{\bm{\Sigma}}\sqrt{{\bm{A}}(t)}{\bm{V}}^{\top}$ (6)
for ${\bm{A}}(t)=\operatorname{diag}(a_{1}(t),\ldots,a_{r}(t))$ with
$a_{i}(t)=\frac{1}{1+(\sigma_{i}e^{2\delta}-1)e^{-2\sigma_{i}t}},~{}i\in[r].$ (7)
Moreover, the time-rescaled factors $a_{i}(\delta t)$ converge to a step function as $\delta\rightarrow\infty$ (limit of vanishing initialization):
$a_{i}(\delta t)\rightarrow\frac{1}{1+\sigma_{i}}\mathds{1}[{t=T_{i}}]+\mathds{1}[{t>T_{i}}],$ (8)
where $T_{i}=1/\sigma_{i}$ and $\mathds{1}[{A}]$ is the indicator function for event $A$. Thus, the $i$-th component is learned at time $T_{i}$ inversely proportional to $\sigma_{i}$.
Proof.
After having set up the analogy of our setting to that of Saxe et al. [2013], Gidel et al. [2019], this is a direct application of [Gidel et al., 2019, Thm. 1]. Specifically, this is made possible in our setting by: (i) interpreting the UFM with square-loss in (5) as a two-layer linear network (ii) recognizing that the covariance of the inputs (which here are standard basis vectors $\bm{e}_{j},j\in\mathbb{R}^{m}$) is the identity matrix, hence the (almost) orthogonality assumption (see [Saxe et al., 2013] and [Gidel et al., 2019, Sec. 4.1]) holds. ∎
The result requires initializing word/context embeddings aligned with the SVD factors of the data-sparsity matrix. While this might appear as a strong assumption, Saxe et al. [2013, 2019] conjectured and verified experimentally that the characterization remains qualitatively accurate under small random initialization. Our experiments with the data-sparsity matrix confirm this - Fig. 18 shows the singular values of the logit matrix during training closely follow the predicted exponential trend in Eq. (7). This reveals that dominant singular factors, corresponding to primary semantic concepts, are learned first. In the limit $t\rightarrow\infty$, the theorem shows convergence to:
${\bm{W}}(t)\rightarrow{\bm{W}}_{\infty}:=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}^{\top}\qquad\text{and}\qquad{\bm{H}}(t)\rightarrow{\bm{H}}_{\infty}:=\mathbf{R}\sqrt{\bm{\Sigma}}{\bm{V}}^{\top},.$ (9)
This aligns with Zhao et al. [2024]’s regularization-path analysis of NTP-UFM with CE loss, where normalized quantities converge as ${\lambda}\rightarrow 0$ (recall Section 3.2):
$\overline{{\bm{W}}}({\lambda})\rightarrow\bm{U}\sqrt{\overline{\bm{\Sigma}}}\mathbf{R}^{\top}\qquad\text{and}\qquad\overline{{\bm{H}}}({\lambda})\rightarrow\mathbf{R}\sqrt{\overline{\bm{\Sigma}}}{\bm{V}}^{\top}.$
Figure 12: (Top) Support matrix and SVD factors of centered support matrix for a synthetic example. (Bottom) Training dynamics of GD minimization of NTP-UFM with square loss (Eq. (5) for two initializations: (i) SVD: initialize ${\bm{W}}$ and ${\bm{H}}$ as per Thm. 1 for $\delta=8$. (ii) Rand: intialize ${\bm{W}}$ and ${\bm{H}}$ random Gaussian scaled to match the norm of SVD initialization. Dynamics with the two initialization are shown in red (SVD) and blue (Rand), respectively. Qualitatively the behavior is similar. Left: Training loss and norms of paramteres. Middle-Left: Convergence of word and context gram-matrices and of logits to the theory predicted by Thm. 1. Middle-Right: Convergence of singular values of logit matrix to those of $\bm{\Sigma}$ (see Thm. 1. Right: Projection of logits to subspace orthogonal to $\bm{U}$ and ${\bm{V}}$; Logits with Rand initialization initially have non-zero projection but it becomes zero as training progresses.
D.1 An example of controlling the rate of learning via reweighting
Consider minimizing the following weighted version of (5):
$\min_{{\bm{W}},{\bm{H}}}~{}~{}\left\|\left(\widetilde{\bm{S}}-{\bm{W}}{\bm{H}}\right)\bm{\Omega}\right\|^{2}\,,$ (10)
where $\bm{\Omega}=\operatorname{diag}([\omega_{1},\ldots,\omega_{m}])$ is a diagonal matrix of weights, one for each context. Here, we consider the STEP-imbalanced one-hot classification setting described in Sec. A.4. Concretely, let the support matrix be
$\bm{S}=\begin{bmatrix}\mathds{I}_{V/2}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}_{V/2\times V/2}\\ \mathbf{0}_{V/2\times RV/2}&\mathds{I}_{V/2}\end{bmatrix}$ (11)
where $R$ is the imbalance ratio and without loss of generality we assumed that the first $V/2$ classes are majorities and that minorities have $1$ example (in our language: context) each. (Thus, the total number of examples is $m=RV/2+V/2=(R+1)V/2$.) Recall that $\widetilde{\bm{S}}=(\mathds{I}_{V}-\frac{1}{V}\mathds{1}_{V}\mathds{1}_{V}^{\top})\bm{S}$.
Recall from Sec. A.4 that $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ where the singular values and the left singular vectors are explicitly given by:
$\bm{\Sigma}=\operatorname{diag}(\begin{bmatrix}\sqrt{R}\mathds{1}_{V/2-1}&\sqrt{(R+1)/2}&\mathds{1}_{V/2-1}\end{bmatrix}),\qquad\bm{U}=\left[\begin{array}[]{ccc}\mathbb{F}&-\sqrt{\frac{1}{V}}\mathds{1}&\mathbf{0}\\ \mathbf{0}&\sqrt{\frac{1}{V}}\mathds{1}&\mathbb{F}\end{array}\right]\in\mathbb{R}^{V\times(V-1)}\,,$
with $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$. According to [Thrampoulidis et al., 2022, Lem. A.3] we also have
${\bm{V}}^{\top}=\left[\begin{array}[]{ccc}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]\in\mathbb{R}^{(V-1)\times m}\,.$
We now set the weight matrix as follows:
[…]
Direct calculation yields
${\bm{V}}^{\top}\bm{\Omega}$ $=\left[\begin{array}[]{ccc}\sqrt{\frac{1}{R}}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{2}{R(R+1)V}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]$ (16)
$=\operatorname{diag}(\begin{bmatrix}\sqrt{\frac{1}{R}}\mathds{1}_{V/2-1}^{\top}&\sqrt{\frac{2}{R+1}}&\mathds{1}_{V/2-1}^{\top}\end{bmatrix})\underbrace{\left[\begin{array}[]{ccc}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{1}{RV}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{1}{V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]}_{=:\widetilde{\bm{V}}^{\top}}=\bm{\Sigma}^{-1}\widetilde{\bm{V}}^{\top}\,.$ (20)
where in the last equation, we recognized the diagonal matrix is equal to $\bm{\Sigma}^{-1}$ and called the other matrix $\widetilde{\bm{V}}\in\mathbb{R}^{m\times(V-1)}$. Thus, the effective sparsity matrix $\widetilde{\bm{S}}_{\bm{\Omega}}=:\widetilde{\bm{S}}\bm{\Omega}$ is equal to
$\widetilde{\bm{S}}_{\bm{\Omega}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}\bm{\Omega}=\bm{U}\widetilde{\bm{V}}^{\top}\,.$
It is now easy to check that $\widetilde{\bm{V}}^{\top}\widetilde{\bm{V}}=\mathds{I}_{V-1}$. Thus, the display above is the SVD factorization of $\widetilde{\bm{S}}_{\bm{\Omega}}$. Note that the singular values of this effective sparsity matrix are all equal, unlike the singular values $\bm{\Sigma}$ of the original data-sparsity matrix $\widetilde{\bm{S}}$. With this at hand, we can show analogous to Proposition 1 that the eigenvalues of ${\bm{W}}(t)$ and ${\bm{H}}(t)$ are now all learned at the same time $T=1$. We also confirm this experimentally in Fig. 13.
Figure 13: (Top) Support matrix and SVD factors of centered support matrix for a one-hot STEP-imbalanced example. (Middle) Training dynamics of GD minimization of NTP-UFM with square loss (Eq. (5); same setup and visualizations as Fig. 12. (Bottom) Training dynamics of GD minimization of NTP-UFM with weighted square loss (Eq. (10)) with weights as in Eq. (12). Note that thanks to the weighting all singular factors are now learned simultaneuously.
Appendix E Extended Discussion on Related Works
Word embeddings and semantic analysis in neural probabilistic language models.
The word2vec architecture Mikolov et al. [2013d, b] and its variants, notably GloVe Pennington et al. [2014], represent seminal early neural probabilistic language models. These simple log-bilinear models, trained on large text corpora, revolutionized word embedding learning. As noted in Zhao et al. [2024], NTP-UFM shares structural similarities with these early models, though in both their work and ours, it serves as a tractable abstraction rather than a practical architecture. Our approach differs by learning both context and word embeddings, following modern practice. The foundational work of Levy and Goldberg [2014] connected word2vec’s geometry to matrix factorization of the pointwise mutual information (PMI) matrix—a specialized word co-occurrence matrix. Subsequent works Levy et al. [2015], Turney and Pantel [2010], Baroni and Lenci [2010] empirically demonstrated semantic interpretations of the PMI matrix’s singular factors and principal components. Building on Zhao et al. [2024], which formalizes a modern version of Levy and Goldberg [2014]’s results, our investigation of concepts differs from this classical literature in two key aspects: (a) We study the NTP setting where both context and word embeddings are learned, yielding concepts that relate to both words and contexts; (b) Our data-sparsity matrix differs fundamentally from classical PMI matrices: it is a centered version of the data support matrix (independent of specific next-token probabilities) and has different structural properties—being orthogonal and non-square, unlike the word2vec setting.
Superposition and feature steering.
Our work was partly motivated by recent compelling literature suggesting that embeddings can be decomposed into linear combinations of a finite set of semantic concepts Bricken et al. [2023], Yun et al. [2023], Park et al. [2023]. These insights from mechanistic interpretability have led to practical applications in "feature steering"—where model behavior can be controlled by manipulating concept representations through addition or subtraction Durmus et al. [2024], Konen et al. [2024]. Our analysis complements the mechanistic interpretability approach by providing a systematic framework for understanding how concepts emerge naturally as principal components from training data statistics. Exploring deeper connections between our theoretical framework and the mechanistic interpretability literature remains an intriguing direction for future work. For completeness, we note that Park et al. [2023, 2024] also investigate geometric properties of concept directions, albeit through fundamentally different technical approaches, assumptions, and perspectives.
Saxe et al.’s closed-form dynamics of two-layer linear network training.
Our work draws inspiration from Saxe et al. [2019]. Conceptually, Saxe et al. [2019] uses a two-layer linear neural network as a theoretical proxy to study the emergence of semantic knowledge in human cognition, providing mathematical justifications for phenomena observed in cognitive semantics literature. A key insight from their work is that even a simple two-layer linear network with orthogonal inputs can yield rich and meaningful conclusions about semantic learning. While two-layer neural networks represent perhaps the simplest instances of non-linear learning, their training dynamics generally remain analytically intractable. However, Saxe et al. [2013] (with aspects later formalized in Gidel et al. [2019]) demonstrated that with square loss, orthogonal inputs, and sufficiently small initialization, these dynamics admit exact closed-form solutions. This mathematical characterization underlies their results on semantic information development through singular factors of the network’s input-output correlation matrix. We make a novel connection to this line of work: the UFM fits perfectly within the framework studied by Saxe et al. [2013, 2019]. Specifically, the UFM can be viewed as a linear two-layer network where the input dimension equals the number of input examples (in our case, the number of contexts $m$). This connection is valuable in two directions: First, the UFM—recently popularized through neural collapse literature (see below)—provides perhaps the most natural and practical setting satisfying Saxe et al.’s seemingly restrictive orthogonal input assumptions. Second, this connection allows us to leverage Saxe et al.’s earlier results in the evolving neural collapse literature. Despite these methodological similarities with Saxe et al. [2019], our work differs in motivation and interpretation. We focus specifically on NTP and how semantic and grammatical concepts emerge from natural language data, rather than general cognitive development. Additionally, we primarily focus on CE loss which is typically used in NTP training.
Neural-collapse geometries.
Our results contribute to the recent literature on the neural collapse (NC) phenomenon Papyan et al. [2020a]. Originally observed in one-hot classification training of DNNs, neural collapse describes two key properties of well-trained, sufficiently expressive DNNs: (1) NC: embeddings of examples from the same class collapse to their class mean, and (2) ETF-geometry: class-mean embeddings form a simplex equiangular tight frame (ETF), being equinorm and maximally separated, with classifier weights exhibiting the same structure and aligning with their respective class-mean embeddings. This phenomenon, consistently observed across diverse datasets and architectures, has sparked extensive research interest, generating hundreds of publications a complete review of which is beyond our scope. Instead, we discuss below the most closely-related line of works. One fundamental direction, which forms the basis for many extensions, focuses on explaining NC’s emergence through the UFM. This model abstracts training as joint optimization of last-layer embeddings (unconstrained by architecture) and classifier weights Mixon et al. [2022], Fang et al. [2021]. Multiple influential works have proven NC emergence by analyzing the UFM’s global optima (e.g., Zhu et al. [2021], Garrod and Keating [2024], Jiang et al. [2023]), with extensions to various loss functions beyond cross-entropy, including square loss Zhou et al. [2022a], Han et al. [2021], Tirer and Bruna [2022], Súkeník et al. [2023, 2024] and supervised contrastive loss Graf et al. [2021], Zhou et al. [2022b], Kini et al. . Most early works maintained the original assumptions from Papyan et al. [2020a]: balanced data (equal examples per class) and embedding dimension $d$ exceeding the number of classes $C$. Recent work has explored $d<C$ settings, though often requiring additional assumptions on the loss function Jiang et al. [2023], Liu et al. [2023]. More substantial progress has emerged in the $d>C$ regime with unbalanced data, where Thrampoulidis et al. [2022] provided a complete characterization for step-imbalanced data (where examples are distributed equally within minority classes and equally within majority classes). They introduced the SELI (simplex-encoding labels interpolation) geometry, showing that logits interpolate a simplex-encoding matrix—a centered version of the one-hot encoding matrix. The embeddings and classifier vectors are then determined, up to rotation and scaling, by the singular vectors of this matrix. The SELI geometry emerges as a special case of the richer geometries characterized in the NTP setting by Zhao et al. [2024]. Together with Li et al. [2023], these works stand alone in extending geometric characterization beyond one-hot encoding—to soft-label and multilabel settings respectively. Specifically, Zhao et al. [2024] analyzes the soft-label setting arising in NTP training on natural language, showing that word and context embeddings are determined by the singular factors of the data sparsity matrix. Our work deepens this understanding by revealing that these SVD factors encode conceptual meaning, thereby extending neural collapse geometry to capture not only the structure of embeddings but also the organization of latent concepts.
Appendix F Additional Experiment Results
F.1 Semantics for Individual Concept
In this section we include the plots illustrating that individual concept do not contain interpretable semantics as discussed in sec 5
Figure 14: Word clouds of top words from individual concept dimensions across datasets, illustrating their lack of human-interpretable semantics without combination. Shown: positive 2nd dimension (Simplified TinyStories), negative 6th dimension (Simplified TinyStories), negative 2nd dimension (Simplified WikiText), positive 4th dimension (Simplified WikiText).
F.2 Experiment on the Semantics of Pretrained Context Embeddings
To identify the semantics in context embeddings from pretrained transformers, we chose GPT-2 Radford et al. [2019] and extracted $50,000$ text sequences, each 11 tokens in length from GPT-2’s training corpus, the OpenWebText dataset. By segmenting these sequences into contexts of varying lengths (1-10 tokens), we generated $500,000$ contexts, which we designate as the "simplified OpenWebText" dataset.
To obtain concepts from pretrained models, instead of performing SVD on $\widetilde{\bm{S}}$ (which is computationally expensive to construct and factorize), we work directly with the embeddings of the pretrained model. For word concepts, we extract ${\bm{W}}$ from GPT-2’s decoder and take the matrix $\bm{U}$ of word concepts to be its left singular matrix. For context concepts, we form a $d\times m$ matrix ${\bm{H}}$ by concatenating the embeddings (last-MLP layer representations before the decoder) of GPT-2 for each context in our subset. This matrix corresponds to only a portion of the “true” embedding matrix for the entire dataset. To ensure the coupling between ${\bm{W}}$ and ${\bm{H}}$ that is inherent to the SVD of $\widetilde{\bm{S}}$ (which we don’t have direct access to), we compute the matrix ${\bm{V}}$ of context concepts as ${\bm{V}}={\bm{H}}^{\top}\mathbf{R}\bm{\Sigma}^{-1/2}$, where $\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ denotes the SVD of ${\bm{W}}$ that we already computed.
We include the orthant-clustering result on GPT2’s pretrained word embedding with $p=4,5,6$. (Fig. 15, 16, 17)
Figure 15: orthant-based clustering on Pretrained word embedding, $p=4$, GPT2
Figure 16: Semantics on Pretrained word embedding, $p=5$, GPT2
Figure 17: orthant-based clustering on Pretrained word embedding, $p=6$, GPT2
F.3 Rate of learning
Fig. 18 shows the evolution of singular values during training for both squared loss and cross-entropy loss as described in Section 4.4.
Figure 18: Evolution of singular values of the logit matrix during training. Both plots show that dominant concepts (corresponding to larger singular values) are learned first. (Left) With squared loss, singular values converge to those of the optimal solution, demonstrating learning saturation. (Right) With CE loss, singular values grow unboundedly while maintaining their relative ordering, reflecting the continuous growth of embedding norms characteristic of CE training.
F.4 $d>V$
Fig. 19 shows the evolution of singular values for squared loss when the network’s embedding dimension is smaller than the vocabulary size, i.e., $d<V$, as described in Section A.1.
Figure 19: evolution of the learned logit matrix’s singular values during training for $3=d<V$, note that the 3 singular values are converging to the $d$ largest singular values
F.5 Semantics from K-means Spectral Clustering on Simplified TinyStories and Wiki-Text
Please see Fig. 20 and 21.
Figure 20: 32-means spectral clustering, $p=5$, Simplified TinyStories
Figure 21: 64-means spectral clustering, $p=6$, Simplified WikiText
F.6 More orthant-based clustering result on Simplified TinyStories and Wiki-Text
We include some examples of orthant-based clustering results in Fig. 22 to 25. We selected $p=4$, $p=5$ for Simplified TinyStories and $p=5$, $p=6$ for Simplified WikiText because we observe the most human-interpretable concepts.
Figure 22: Orthant-based clustering, $p=4$, Simplified TinyStories
Figure 23: Orthant-based clustering, $p=5$, Simplified TinyStories
Figure 24: Orthant-based clustering, $p=5$, Simplified Wiki-Text
Figure 25: Orthant-based clustering, $p=6$, Simplified Wiki-Text

### v2

@@
-On the Geometry of Semantics in Next-token Prediction
+Geometry of Semantics in Next-Token Prediction: How Optimization Implicitly Organizes Linguistic Representations
 Abstract
-Modern language models demonstrate a remarkable ability to capture linguistic meaning despite being trained solely through next-token prediction (NTP).
-We investigate how this conceptually simple training objective leads models to extract and encode latent semantic and grammatical concepts.
-Our analysis reveals that NTP optimization implicitly guides models to encode concepts via singular value decomposition (SVD) factors of a centered data-sparsity matrix that captures next-word co-occurrence patterns.
-While the model never explicitly constructs this matrix, learned word and context embeddings effectively factor it to capture linguistic structure.
-We find that the most important SVD factors are learned first during training, motivating using spectral clustering of embeddings to identify human-interpretable semantics, including both classical k-means and a new orthant-based method directly motivated by our interpretation of concepts.
-Overall, our work bridges distributional semantics, neural collapse geometry, and neural network training dynamics, providing insights into how NTP’s implicit biases shape the emergence of meaning representations in language models.
+We investigate how next-token prediction (NTP) optimization leads language models to extract and organize semantic structure from text.
+Our analysis, based on a tractable mathematical model and controlled synthetic data, reveals that NTP implicitly guides models to factor a centered support matrix encoding context-to-next-token co-occurrence patterns via singular value decomposition (SVD).
+While models never explicitly construct this matrix, learned word and context embeddings converge to its SVD factors, with singular vectors encoding latent semantic concepts through their sign patterns.
+We demonstrate that concepts corresponding to larger singular values are learned earlier during training, yielding a natural semantic hierarchy where broad categories emerge before fine-grained ones.
+This insight motivates orthant-based clustering, a method that combines concept signs to identify interpretable semantic categories.
+We validate our findings on synthetic datasets and pretrained language models, recovering diverse semantic structures such as grammatical categories, named entity types, and topical distinctions (medical, entertainment).
+Our work bridges classical distributional semantics and neural collapse geometry, characterizing how gradient-based optimization implicitly determines both the matrix representation and factorization method that encode semantic structure.
 1 Introduction
-Next-token prediction (NTP) is conceptually simple: predict the next word given a preceding sequence (aka context). Yet it trains models with remarkable ability to capture meaning.
-A foundational step toward understanding how NTP leads to models acquiring semantic information is to first understand how explicit textual inputs—words and contexts—are represented in the models’ $d$-dimensional word and context vector representations (embeddings).
-Zhao et al.
-[2024] recently proved that in well-trained models, the geometry of these learned representations is characterized by the factorization of a data-sparsity matrix $\widetilde{\bm{S}}$ whose entries encode whether a particular word follows a given context in the training corpus.
-While this characterization explains how models encode explicit training signals, it does not address the deeper question about semantic learning.
-Language conveys meaning through latent concepts that exist beyond explicit (context, word) pairs, leading us to investigate: How do NTP-trained models extract and encode these latent concepts?
-For instance, how do they capture semantic dichotomies like male/female, or grammatical categories like nouns and verbs?
-In other words, how does latent information shape the embedding geometry, and does this suggest ways to identify interpretable semantic structures in the embedding space?
-These questions are challenging because linguistic concepts are never explicitly given as inputs in the NTP objective.
-Yet, by analyzing the geometry of learned representations and its relation to the data-sparsity matrix, we show that latent linguistic structure emerges from the latter.
-Concretely, our contributions are summarized as follows.
-Geometric emergence of latent concepts: We demonstrate that latent linguistic concepts learned by a large neural model through NTP emerge as principal components in the singular value decomposition of a centered data-sparsity matrix $\tilde{\mathbf{S}}$ (defn.
-in Sec.
-3.1, see Fig.
-1 for illustration).
-While $\widetilde{\bm{S}}$ is never explicitly formed during training, its structure naturally shapes the geometry of learned representations: concepts can be recovered as weighted combinations of word and context embeddings, where the weights are determined by the singular vectors of $\tilde{\mathbf{S}}$.
-This extends the characterization of the unconstrained-features model (UFM) by Zhao et al.
-[2024], revealing how embeddings inherently organize around latent concept dimensions.
-Rate of learning: Singular values of the data-sparsity matrix quantify the significance of their corresponding concepts.
-By recognizing a connection between UFM with square-loss and the closed-form training dynamics of two-layer neural networks with orthogonal inputs derived by Saxe et al.
-[2013], we show that concepts associated with larger singular values are learned faster during training.
-Semantics [^1] We broadly define “semantic” to encompass linguistic factors that humans consider when producing text, including grammatical (e.g., tense), syntactic (e.g., verb), and semantic (e.g., sentiment) elements.
-as orthant-based clustering: Not all concepts have human-interpretable meaning.
-Guided by the insight that top concepts are learned first during training, we find that human-interpretable semantic information is captured through specific positive/negative configurations (indicating how present/absent the specific concept is in that word or context) across a few top singular concept dimensions (see Fig.
-2B for illustration).
-Identifying semantics through such sign configurations geometrically manifests as (spectral) clustering of word/context embeddings in orthant slices.
-Through experiments, we verify that this approach identifies distinct human-interpretable linguistic categories.
-(A) Left Singular Matrix ($\bm{U}$)Refer to caption(B) Singular ValuesRefer to caption(C) Right Singular Matrix (${\bm{V}}$)Refer to caption
-Figure 1: Illustration of semantic information in SVD of $\widetilde{\bm{S}}$ (rank=6).
-(A) Left singular matrix (U) where rows are analyzer vectors and columns are singular dimensions.
-The first dimension distinguishes between plants (“apple”, “tulip”, negative values) and animals (“dog”, “shark”, positive values).
-The second dimension further differentiates mammals (“dog”, “cat”, positive) from other animals (“trout”, “eagle”, negative).
-(B) Singular values in descending order of importance.
-(C) Right singular matrix (V) where the first dimension separates plant properties (“uses sunlight for food”) from animal properties (“hunts in groups”).
-The second dimension appears more complex: for animal properties, it contrasts mammal-specific traits (“stays warm-blooded”, positive) against aquatic traits (“lives in water”, negative), while for plant properties, it differentiates growth characteristics (“grows roots”, positive) from reproductive features (“reproduces with pollen”, negative).
+Models trained with next-token prediction (NTP) show a remarkable ability to capture and represent linguistic meaning.
+How does this semantic ability emerge during training?
+A primary challenge in addressing this question lies in quantifying “meaning”.
+We therefore focus specifically on understanding the semantic organization of learned representations and the mechanisms by which it arises.
+Our goal is fundamentally different from the substantial recent interest in post-hoc evaluation of interpretable semantics in Large Language Models (LLMs).
+Techniques like dictionary learning and sparse autoencoders for model steering Cunningham et al.
+[2023], Bricken et al.
+[2023], Turner et al.
+[2023], Li et al.
+[2024]—tracing back to similar work with static word embeddings for word2vec architectures Murphy et al.
+[2012], Faruqui et al.
+[2015], Arora et al.
+[2018], Zhang et al.
+[2019]—aim to extract interpretable features from trained models.
+In contrast, we seek to understand the specific mechanisms by which gradient-descent-based training, in conjunction with the structure of textual input data, guides a language model to organize its representations into semantically interpretable structures.
+Our starting point is the recent work by Zhao et al.
+[2024], who characterized the geometry of the model’s explicit textual input representations, i.e., $d$-dimensional word/context embeddings.
+At convergence, this geometry is precisely characterized by the factorization of a specific data-sparsity matrix $\widetilde{\bm{S}}$, which encodes co-occurrence patterns in the training corpus: its entries indicate whether a particular word follows a given context.
+This characterization, however, leaves open the question: How do these textual input patterns impact the geometry of implicit linguistic information, such as grammatical or semantic categories, through the optimization of the NTP objective?
+Specifically, we ask three subquestions:
+Since semantic information is not explicitly labeled in the data,
+Q1. what latent signals implicitly guide NTP optimization to specific structures?
+Q2. How are these structures related to human-interpretable semantics?
+Q3. How fast do these semantic structures emerge during training?
+Figure 1: Illustration of C1-3.
+Details in Sec.
+3.
+(A) SVD of the data-sparsity matrix $\widetilde{\bm{S}}$ (Eq.
+(2)), derived from the training data in Fig.
+2(C).
+Columns of $\bm{U}$ and $\bm{V}$}, serve as word and context concept basis vectors, respectively: Their entries encode an association score with sign indicating the direction (association/opposition) between word/contexts and concepts.
+The singular values indicate the strength of each concept.
+(B) Combining a few top concepts according to the specific sign patterns of concept basis, which we call orthant-based clustering, reveals human-interpretable semantics.
+(C) Gradient-descent-based NTP optimization leads to concepts corresponding to larger singular values being learned first.
+This corresponds to the emergence of coarse semantics before fine-grained ones, which we empirically validate with visualizations of the trained model’s output confusion matrices.
+Towards addressing these questions, we adopt a bottom-up, theory-guided approach.
+We build our insights by analyzing the optimization of an abstract mathematical model, which isolates the essential components of NTP optimization in a tractable framework (Sec.
+2).
+To validate and further shape these insights, we carefully construct synthetic data settings and formally verify our findings, which can otherwise be difficult to precisely quantify in complex, real-world datasets (Sec.
+3).
+Finally, we demonstrate that the insights gained extend qualitatively to real data and models (Sec.
+4).
+Concretely, our contributions are:
+C1.
+Interpreting latent signals (Sec.
+3.1): We show that the singular vectors of $\tilde{\mathbf{S}}$ encode latent concepts: the sign of each entry indicates whether a word or context supports or opposes a concept, and the magnitude reflects the strength of association.
+This equips the abstract geometry in Zhao et al.
+[2024] with semantic structure.
+Fig.
+1A visualizes semantics aligning with concept signs.
+C2.
+Establishing human-interpretable semantic structure (Sec.
+3.2): Building on C1, we introduce an orthant-based clustering method that identifies interpretable semantic categories by combining a few top concepts according to their sign patterns.
+We empirically demonstrate that this method recovers a wide array of interpretable linguistic structures—such as part of speech, verb tense, and named entity types—on both synthetic data and real datasets like TinyStories and WikiText-2, where the data-sparsity matrix is explicitly computable.
+Furthermore, we show in GPT-2, BERT, and Qwen that increasing the number of singular components systematically reveals finer-grained semantics.
+Fig.
+1B visualizes a hierarchy of interpretable semantics via orthant-based clustering.
+C3.
+Analyzing semantic emergence (Sec.
+3.3): Beyond convergence, we analyze semantic emergence during training.
+Theoretically, we link NTP-UFM gradient descent dynamics to the framework of Saxe et al.
+[2013], showing that top concepts, those corresponding to larger singular values, are learned earlier.
+Fig.
+1C visualizes this phenomenon: broad semantics corresponding to top concepts are formed early during training.
+These contributions bridge classical distributional semantics (e.g., Ding et al.
+[2006], Levy et al.
+[2015]) and recent neural collapse (NC) geometries Papyan et al.
+[2020a], Zhao et al.
+[2024].
+By employing tractable abstractions from the latter, we provide theoretical grounding for the former: we show that gradient-based optimization implicitly determines both the matrix representation (data-sparsity matrix) and factorization method (SVD) for encoding semantics, rather than requiring hand-engineered choices (Sec.
+5).
+Conversely, we extend the scope of NC geometries beyond simple classification settings to characterize the underlying hierarchical semantic representations.
+Notations.
+For any integer $k$, $k]:=\{1,\ldots,k\}$.
+Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$, respectively.
+For matrix $\bm{A}$, $\bm{A}[i,j]$ is its $(i,j)$-th entry, and for vector $\bm{a}$, $\bm{a}[i]$ is its $i$-th entry.
+$\mathds{1}$ is the all-ones vector and $\mathds{I}$ is the identity matrix.
+2 Background: Geometry of Word and Context Representations
+The NTP objective minimizes the cross-entropy (CE) loss between the model’s output distribution given a context—a sequence $\bm{z}$ of tokens from vocabulary of size VV—and the one-hot representation of the next-word/token.
+Since multiple tokens can follow a given distinct context $\bm{z}_{j}$ in a dataset, this is equivalent to minimizing CE between the model’s output distribution and an associated soft-label vector $\hat{\bm{p}}_{j}$ representing the conditional frequency distribution of next tokens for that specific context.
+Because not all vocabulary tokens are valid next-tokens for a given context (even when ignoring the finite size of the data), the soft-label vectors $\hat{\bm{p}}_{j}$ are sparse.
+For a dataset with $m$ distinct contexts, define the support matrix $\bm{S}\in\{0,1\}^{V\times m}$ of the soft-label matrix ${\bm{P}}=[\hat{\bm{p}}_{1},\ldots,\hat{\bm{p}}_{m}]\in\mathbb{R}^{V\times m}$, such that $\bm{S}[z,j]=1\,\Leftrightarrow\,{\bm{P}}[z,j]>0$.
+Building on this formulation, Zhao et al.
+[2024] investigated how the soft-label structure shapes the learned geometry when optimizing the NTP objective with gradient descent.
+Specifically, they studied word embeddings (the rows of the $V\times d$ decoding matrix $\bm{W}$}) and context embeddings (the deep $d$-dimensional representations $\{h(\bm{z}_{j})\}_{j\in[m]}$ of distinct contexts) of a model at convergence.
+For tractability, they relied on two modeling simplifications: (1) using the limit of vanishing ridge regularization as a proxy for the convergence limit of gradient-descent iterations, and (2) assuming a sufficiently expressive neural network that optimizes context embeddings freely, instead of abiding by an architecture-specific parameterization.
+The first simplification is standard in the implicit-bias optimization literature Ji et al.
+[2020], Wei et al.
+[2019], Tarzanagh et al.
+[2023], while the second has been used for language models by Yang et al.
+[2017] and recently popularized in the neural collapse literature Mixon et al.
+[2022], Wojtowytsch [2021], Fang et al.
+[2021].
+Taken together, these abstractions yield a tractable model for NTP optimization, which we refer to as the unconstrained features model (UFM) for NTP training, or NTP-UFM in short: [^1] Note the structural similarity of this log-bilinear objective to word2vec architectures Mikolov et al.
+[2013c], Pennington et al.
+[2014].
+However, NTP-UFM serves as a tractable analytical abstraction rather than a direct practical architecture.
+$\min_{{\bm{W}},{\bm{H}}}\mathcal{L}({\bm{W}}{\bm{H}};{\bm{P}})+{{\lambda}}\|{\bm{W}}\|^{2}+{{\lambda}}\|{\bm{H}}\|^{2}\,.$ (1)
+NTP-UFM jointly optimizes the matrices of word embeddings ${\bm{W}}\in\mathbb{R}^{V\times d}$ and context embeddings ${\bm{H}}:=[\bm{h}_{1},\ldots,\bm{h}_{m}]\in\mathbb{R}^{d\times m}$.
+Here, $\bm{h}_{j}$ represents $h(\bm{z}_{j})$ and is freely optimized for each context; $\lambda$} is the regularization parameter; and $\mathcal{L}$ is the CE loss acting on the logit matrix $\bm{L}={\bm{W}}{\bm{H}}$, parameterized by the soft-label matrix $\bm{P}$}.
+Analyzing NTP-UFM, for large embedding dimensions $d\geq V$,  [^2] This does not restrict the number of contexts ($m$, potentially $\gg d$), enabling the study of rich semantic catgories (Sec.
+3).
+Also, qualitative insights extend to real models (Sec.
+4.2).
+and $\lambda\rightarrow 0$, Zhao et al.
+[2024] show that:
+1.
+Logits: The logit matrix $\bm{L}$ grows unboundedly during training.
+After normalization, it converges in direction to a matrix $\bm{L}^{\rm{mm}}$ that can be explicitly computed given the data support matrix $\bm{S}$.
+2.
+Word/Context Embeddings: Like logits, word and context embeddings grow unboundedly in magnitude, but converge in direction to $\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$ and $\mathbf{R}^{\top}\sqrt{\bm{\Sigma}}{\bm{V}}^{\top}$, respectively.
+Here, $\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ is the SVD of $\bm{L}^{\rm{mm}}$ and $\mathbf{R}$ is a partial orthogonal matrix.
+3. Data-sparsity matrix as proxy: An accurate proxy for $\bm{L}^{\rm{mm}}$ that is very easy to compute is the centered support matrix
+$\widetilde{\bm{S}}:=(\mathds{I}_{V}-\nicefrac{{1}}{{V}}\cdot\mathds{1}_{V}\mathds{1}_{V}^{\top})\,\bm{S}\,$ (2)
+$\widetilde{\bm{S}}[z,j]=\begin{cases}1-\frac{1}{V}\sum_{z\in[V]}\bm{S}[z,j]&\text{if $\bm{S}[z,j]=1$}\\ -\frac{1}{V}\sum_{z\in[V]}\bm{S}[z,j]&\text{if $\bm{S}[z,j]=0$}\end{cases}\,$,
+which we refer to as the data-sparsity matrix. Note that unlike the support matrix $\bm{S}$, the entries of $\widetilde{\bm{S}}$ are not binary due to the centering operation.
+Taken together, Zhao et al.
+[2024] establishes how textual input structure maps through NTP optimization to word and context representation geometry.
+Textual input structure is summarized through the SVD of the data-sparsity matrix, where $\bm{U}\in\mathbb{R}^{V\times r}$, ${\bm{V}}\in\mathbb{R}^{m\times r}$ with $\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}$, singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered as $\sigma_{1}\geq\ldots\geq\sigma_{r}>0$, and rank $r\leq V-1$:
+$\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}.$ (3)
+The geometry of learned representations is characterized by word and context embeddings that converge in the direction to $\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$ and $\mathbf{R}^{\top}\sqrt{\bm{\Sigma}}{\bm{V}}^{\top}$, respectively, with $\mathbf{R}$ a partial orthogonal matrix.
 Refer to caption(A) Embedding geometryRefer to caption(B) Orthant-based clusteringRefer to caption(C) Training data
-Figure 2: NTP-trained word embeddings naturally organize into semantic clusters.
-(C) Training data (showing ${\bm{P}}^{\top}$) consists of pairs using the template “The organism that [attribute] is [subject]”, where attributes are rows and subjects are columns.
-(A) When trained with $d>V>rank(\widetilde{\bm{S}})=6$ and visualized in 3D PCA projection, both words (dots) and contexts (triangular marks) cluster by semantic category—animals (green/blue), and plants (red/orange).
-This demonstrates how semantic structure emerges naturally from the underlying data sparsity pattern, a process our work formalizes mathematically.
-(B) Visualization of orthant-based clustering and hierarchical semantic structure in a 3D subspace of context-analyzer vectors.
-Selected dimensions $\mathcal{K}=\{k_{1},k_{2},k_{3}\}={1,2,4}$ correspond to columns of ${\bm{V}}$ in Fig.
-1C.
-Colored blocks represent different sign configurations: Green ($p=1$): $c_{k_{1}}=-1$ captures "animal" semantic.
-Blue ($p=2$): $[c_{k_{1}},c_{k_{2}}]=[-1,-1]$ refines to "mammal" semantic.
-Yellow ($p=3$): $[c_{k_{1}},c_{k_{2}},c_{k_{3}}]=[-1,-1,-1]$ further specifies "feline" semantic.
-This nested structure illustrates how increasing the number of active dimensions ($p$) yields progressively finer-grained semantic categories, where "mammal" is a subset of "animal", and "feline" is more specific within "mammal".
-Contexts shown with same color within each suborthant are example members according to Defn.
+Figure 2: Optimizing NTP on (context,next-word) input pairs yields a specific geometry of representation of those inputs; orthant-based clustering reveals how this geometry is organized on the basis of interpretable semantic categories.
+(A) Empirical 3D PCA visualization of learned word(dots)/context(triangular marks) embeddings from two-layer transformer trained on the dataset in (C) reveals a clear clustering by semantic category (plants in green/blue, animals in red/orange), despite no explicit semantic signal at training.
+(B) We show (Sec.
+3.2) that semantic categories are organized into a hierarchy of sub-orthants defined by concepts combined with respect to their sign patters.
+For example, Felines occupy a 2D sub-orthant (combining the second and fourth concepts), while Mammals and Animals occupy progressively larger sub-orthants.
+(C) The sparse training data matrix (rows = attributes, columns = subjects) follows the template: “The organism that [attribute] is [subject].”
+3 Geometry of Semantics
+The previous section establishes that word and context embeddings converge to the SVD factors of the data-sparsity matrix $\widetilde{\bm{S}}$.
+But: (i) What latent signals implicitly guide optimization to produce semantic structures within this geometry?
+(ii) How are these related to human-interpretable semantics?
+(iii) How fast do they emerge during training?
+To motivate our analysis, consider Fig.
+2A visualizing word/context embeddings of a transformer trained on (sequence,next-token) pairs from the sparse soft-label matrix $\bm{P}$} shown in Fig.
+2C.
+Despite no explicit semantic supervision during training, clear distinctions emerge.
+For example, animal-related words and contexts cluster on the left, while plant-related ones cluster on the right.
+What drives such an implicit semantic structure?
+3.1 Identifying Concepts as Latent Signals
+Our key insight is to examine the structure of the SVD factors of the data-sparsity matrix $\widetilde{\bm{S}}$.
+Recall word embeddings take the form ${\bm{W}}=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$, where $\bm{U}$ encodes the row-space of $\widetilde{\bm{S}}$.
+Since rows of $\widetilde{\bm{S}}$ (and, thus, of $\bm{U}$) correspond to different words, their inner product captures word similarity: words appearing in similar contexts have high similarity.
+However, semantic relationships extend beyond simple pairwise similarity or synonymy.
+To reveal such semantic structure, we instead focus on the columns of $\bm{U}$, which correspond to $r$ latent features in the representation space.
+We term each SVD dimension $k\in[r]$ a concept.
+Adopting terminology from Saxe et al.
+[2019], we call the columns $\mathbf{u}_{k}\in\mathbb{R}^{V}$ and $\mathbf{v}_{k}\in\mathbb{R}^{m}$ of matrices $\mathbf{U}$ and $\mathbf{V}$ the "word analyzer vectors" and "context analyzer vectors," respectively.
+These vectors represent the alignment of words and contexts with each concept: for a word $z$ and concept $k$, the sign of $\mathbf{u}_{k}[z]$ indicates association ($\mathbf{u}_{k}[z]>0$) or opposition ($\mathbf{u}_{k}[z]<0$) to the concept, while its magnitude quantifies the strength of this relationship.
+Words for which $|\bm{u}_{k}[z]|$ is small are neutral to this concept.
+The same interpretation applies to context components $\mathbf{v}_{k}[j]$.
+We can explicitly represent concepts in the embedding space by defining $d$-dimensional concept representations $\bm{u}^{\rm{d}}_{k}={\bm{W}}^{\top}\bm{u}_{k}$ and ${\bm{v}}^{\rm{d}}_{k}={\bm{H}}{\bm{v}}_{k}$ for $k\in[r]$ as projections onto the word and context spaces, respectively.
+These can also be viewed as weighted averages of word or context embeddings (similar to classical semantic axes) where the entries of $\bm{u}_{k}$ (or ${\bm{v}}_{k}$) reflect each word’s (or context’s) relevance to the concept.
+For an initial confirmation of this interpretation of SVD factors, we revisit the dataset with syntax "The organism that [attribute] is [subject]", [^3] For visual clarity, this design isolates semantic concept extraction by minimizing grammatical influences.
+shown in Fig.
+2C.
+Fig.
+1A shows the columns of $\bm{U}$ and $\bm{V}$}, that is, the word and context analyzer vectors.
+Inspecting them reveals how they can capture semantic information through their sign patterns: For example, the first column/dimension is such that plant-related words and contexts have positive entries, while animal-related ones have negative entries, indicating that this concept represents the semantic distinction between plants and animals.
+However, the figure also reveals that not all individual concepts correspond to human-interpretable categories.
+Instead, we posit that interpretable semantics emerge from specific combinations of several concepts.
+We formalize this in the next section and demonstrate in Sec.
+4 that it yields interpretable semantics on larger-scale datasets.
+3.2 Interpretable Semantics via Orthant-based Clustering
+We extend the sign-based interpretation of individual concepts to combinations of multiple concepts.
+Our key claim is that combining a few top concepts according to their sign patterns, a method we call orthant-based clustering, yields interpretable semantic categories.
+For a selected set $\mathcal{K}=\{k_{1},\ldots,k_{p}\}$ of $p\leq r$ concepts, there are $2^{p}$ possible sign configurations $C=[c_{k_{1}},\ldots,c_{k_{p}}]\in\{\pm 1\}^{p}$, where $c_{k_{i}}\in\{\pm 1\}$ indicates whether dimension $k_{i}$ contributes positively or negatively to a semantic category.
+Each configuration defines a potential semantic category with member words and their degrees of association with that category, formalized as follows:
+Definition 1.
+For $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ where each $c_{k_{i}}=\pm 1$ for concept dimensions $k_{i}\in[r]$:
+$\bullet$ A word $z$ is a member of $C$ if and only if $\texttt{sign}(\bm{u}_{k_{i}}[z])=c_{k_{i}}$ for all $k_{i},i\in[p]$.
+$\bullet$ The typicality of a member $z$ of $C$ is defined as: ${\rm{Typicality}}(z;C)=\sum\nolimits_{i\in[p]}\big|\bm{u}_{k_{i}}[z]\big|\,.$
+Analogous definitions apply to contexts.
+Geometrically, configuration $C$ corresponds to an orthant in the $p$-dimensional subspace spanned by selected concepts.
+Membership indicates which word embeddings lie within this orthant, while typicality measures the $\ell_{1}$-norm of a word’s projection onto the selected concept dimensions.
+[…]
+This reveals a natural hierarchical structure in orthant-based clustering, where the number of concept dimensions determines semantic granularity: broad categories (e.g., verbs) require fewer dimensions, while specific categories (e.g., past-tense verbs) need more dimensions.
+See also Sec.
+4.
+3.3 Semantic Emergence
+We now show that the semantics’ hierarchical structure is reflected in their emergence during training: broad categories are learned earlier than fine-grained ones.
+We formalize this by tracking the rate at which concepts are acquired throughout training.
+We find that concepts aligned with dominant singular directions (those corresponding to larger singular values) are learned first.
+Our analytic NTP-UFM model quantifies this claim as follows.
+Theorem 1.
+Consider gradient-descent minimization of NTP-UFM (Eq.
+(1)) with square loss.
+Assume infinitesimal step-size and spectral initialization ${\bm{W}}(0)=e^{-\delta}\bm{U}\mathbf{R}^{\top},{\bm{H}}(0)=e^{-\delta}\mathbf{R}{\bm{V}}^{\top}$ for partial orthogonal matrix $\mathbf{R}\in\mathbb{R}^{d\times r}$ ($\mathbf{R}^{\top}\mathbf{R}=\mathds{I}_{r}$) and initialization scale $e^{-\delta}$.
+Then, word and context embeddings at iteration $t$ evolve as ${\bm{W}}(t)=\bm{U}\sqrt{\bm{\Sigma}}\sqrt{{\bm{A}}(t)}\mathbf{R}^{\top}$ and ${\bm{H}}(t)=\mathbf{R}\sqrt{\bm{\Sigma}}\sqrt{{\bm{A}}(t)}{\bm{V}}^{\top}$ for ${\bm{A}}(t)=\operatorname{diag}(a_{1}(t),\ldots,a_{r}(t))$ with
+$a_{i}(t)=\frac{1}{1+(\sigma_{i}e^{2\delta}-1)e^{-2\sigma_{i}t}},~i\in[r].$ (4)
+Therefore, the $i$-th concept converges ($a_{i}(t)\rightarrow 1$) with exponential rate of $2\sigma_{i}$: concepts corresponding to larger singular values are learned first.
+Proof and justification of the spectral initialization are in App. C.2. Our key insight is bridging the UFM with Saxe et al. [2013]’s study of closed-form learning-dynamics of two-layer linear nets.
+To empirically verify this in natural language, we track concept learning through confusion matrices at different training timestamps.
+Since NTP performs soft-labeled classification where contexts can have multiple valid next-tokens, we define "effective classes" as the distinct sets of possible next-tokens (i.e., support sets) shared by at least one context in the dataset.
+For the dataset Fig.
+2C, there are 9 such classes.
+Fig.
+1C shows $9\times 9$ confusion matrices at four training timestamps from a 2-layer transformer; entry $(c,c^{\prime})$ indicates how many contexts from class $c$ are predicted to belong to class $c^{\prime}$.
+The diagonal entries show the number of correct classifications, with the total number of contexts per class shown after the slash.
+We observe that certain effective classes naturally correspond to semantic categories, as contexts sharing semantic meaning (e.g., describing animal attributes) tend to share the same support set (e.g., animal names).
+We annotate these associations with labels ("Animals," "Plants") in brackets on the confusion matrix axes.
+The visualization reveals that the model first learns classes with coarse semantic associations (e.g., "Animals" vs.
+"Plants" at step 20), then progressively distinguishes classes with finer-grained semantic associations (e.g., "Lives in water (fish)" by step 40).
+This progression reflects that coarse semantic structures, corresponding to concepts with larger singular values, are learned earlier in training than finer-grained ones that require combining more concept dimensions.
+To isolate the essential mechanism driving this progressive emergence, we identify the minimal data structure required: class imbalance alone suffices.
+Fig.
+10 in Appendix demonstrates this on imbalanced MNIST, where the model learns distinctions in order of their corresponding singular values.
+Analyzer vectors with larger singular values encode majority-majority distinctions (learned first), the middle singular value encodes the majority-minority distinction (learned next), and smaller singular values encode minority-minority distinctions (learned last).
+This confirms that the progressive emergence of concept dimensions, ordered by singular value magnitude, stems from the fundamental optimization geometry determined by $\widetilde{\bm{S}}$ rather than being specific to linguistic structure.
+However, while even minimal imbalance produces meaningful categorical distinctions, richer sparsity patterns in $\bm{S}$, as naturally arise in the soft-labeled language setting, enable progressively richer semantic concepts to emerge from its SVD, as we show in Sec.
+4.
+4 Experimental Validation
+Datasets.
+We construct datasets by subsampling from two distinct corpora: TinyStories [Eldan and Li, 2023] and WikiText-2 [Merity et al., 2017].
+For both, we fix vocabulary size V=1V=1k and extract m=10m=10k contexts by sampling frequent sequences of lengths 2–6 from the original corpora, creating Simplified TinyStories and Simplified WikiText.
+Models.
+To find semantics in models trained on controlled data, we train decoder-only transformers of varying sizes (6 or 12 layers, 1024 hidden dimensions) on Simplified datasets using AdamW until convergence (see Appendix D.2) for training details).
+We also analyze pretrained models starting with GPT-2-medium.
+To validate findings across architectures, scaling, and multi-language settings, we also analyze Qwen-7B (7B parameters with 4096-dimensional embeddings) [Bai et al., 2023].
+Finally, while we focus primarily on NTP loss, the underlying structure of context-to-next-token co-occurrence is also present in masked language modeling (MLM) loss, where the context consists of surrounding tokens and the model is trained to predict randomly masked tokens within the sequence.
+To verify our analysis applies to here as well, we experiment with BERT-base [Devlin et al., 2018].
+4.1 Methodology
+Our approach differs depending on whether we have access to the training corpus.
+Simplified datasets: We explicitly construct $\widetilde{\bm{S}}$ and compute word and context analyzer vectors for each concept via SVD (Sec.
+3.1).
+Pretrained models: We do not have access to their pretraining corpora, making it infeasible to construct $\widetilde{\bm{S}}$ and compute its SVD.
+We thus resort to an approximation: we extract the decoder matrix $\bm{W}$}, then use Arnoldi decomposition to compute their left singular vectors $\bm{U}$, which serve as proxies for word analyzer vectors.
+For both cases, once we have analyzer vectors, we apply orthant-based clustering (Sec.
+3.2).
+We combine the top-$p$ concepts $\{1,\ldots,p\}$ in successive order, starting from the most dominant.
+We find that choosing $p$ as small as 5 to 7 and inspecting word or context membership across different sign configurations (i.e., orthants) typically suffices to reveal interpretable semantic patterns.
+To interpret semantics, we visualize top 40 words/contexts associated with each orthant.
+“Association” and “top” are evaluated per Defn.
 1.
-Connecting distributional semantics and neural collapse (NC) geometries: On the one hand, linking to classical distributional semantics approaches based on heuristic forms of matrix factorization of data co-occurrences, we find using model abstractions popularized in the NC literature that NTP implicitly favors SVD factorization of the centered data-sparsity matrix.
-On the other hand, we broaden the scope of NC geometries of embeddings to encompass geometries of latent semantics – making the connection explicit, we show the emergence of semantics even for the default one-hot data setting in the NC literature.
+We use word clouds with size proportional to typicality score.
+Symbol # indicates token start.
+4.2 Results
+Individual concepts are not interpretable.
+We verify that not all individual concepts correspond to linguistically interpretable categories (Sec.
+3.1).
+Fig.
+12 in App.
+demonstrates this for Simplified TinyStories and WikiText.
+This aligns with findings in the recent literature [Chersoni et al., 2021, Piantadosi et al., 2024, Elhage et al., 2022].
+Past-tense verb
+Present-tense verb
+Preposition
+Proper names
+Month
+Number
+Unit
+Preposition/Conjunction
+Figure 3: Semantics from orthant-based clustering on $\widetilde{\bm{S}}$ in Simplified TinyStories (Top) and Simplified Wikitext (Bottom).
+Semantic interpretations and combination configurations are shown in the title.
+Positive
+Negative
+Chemical
+Medical
+Cooperation
+Digital entertainment
+Cultural
+Societal
+Figure 4: Semantics identified by orthant-based clustering on GPT-2’s word embeddings.
+Combining concepts with sign configurations reveals semantics.
+We verify that combining (non-interpretable) individual concepts reveals interpretable semantics via orthant-based clustering (Sec.
+3.2).
+Fig.
+3 shows human-interpretable semantic categories identified by orthant-based clustering in Simplified TinyStories and WikiText.
+In Fig.
+11 in Appendix, we further demonstrate the hierarchical semantic structure that emerges as we combine more concepts, illustrated through verb forms organized by grammatical properties.
+Starting with only the top concept ($k=1$), the categories remain broad and difficult to interpret.
+However, as we incorporate additional concepts, increasingly refined semantic categories emerge, ultimately distinguishing between different verb forms (past tense, present tense, and present continuous).
+Orthant-based clustering can be viewed as instance of spectral clustering Von Luxburg [2007].
+We thus also explore k-means spectral clustering over concept subspaces in App.
+D.8.
+While this also reveals interpretable categories, orthant-based clustering offers advantages: (1) It provides a direct link between optimization and categories by explicitly interpreting sign patterns of SVD dimensions as semantic indicators.
+(2) It has built-in selectivity: not all $2^{p}$ potential orthants yield semantically coherent groupings—only a meaningful subset does.
+(3) This control over $p$ enables semantic interpretation of membership and typicality, and allows exploration of different semantic granularities.
+Figure 5: Semantics from transformers of different depth, both trained on Simplified TinyStories.
+Figure 6: Context/word semantics share identical configurations.
+Left and right pairs illustrate the “verb+to infinitive” and “was so/very + adj” structures, confirming semantic alignment between contexts and their in-support words.
+Relation of words and context semantics.
+Recall from Secs.
+3.1,3.2 that semantic clustering of both word and context embeddings is guided by the SVD of the same data-sparsity matrix $\widetilde{\bm{S}}$.
+This coupling predicts that words and contexts sharing semantic properties should cluster together in embedding space.
+Fig.
+2A illustrated this phenomenon in a simple setting, where plant-related words and contexts clustered separately from animal-related ones.
+We now demonstrate this in more complex datasets: We apply orthant-based clustering to words and contexts using the same signature pattern across a subset of concept dimensions extracted from Simplified TinyStories.
+Fig.
+6 shows that contexts with modals like “want/like/love/have to” align with clusters of infinitive verbs; intensifiers like “was so/very” correspond to descriptive adjectives.
+This duality reflects how contexts encode meaning both via their surface words and via the semantics of likely next tokens—a view naturally supported by the predictive nature of NTP.
+Semantics in pretrained models.
+We find that simplified datasets primarily reveal grammatical/syntactic structure due to limited size and context diversity.
+To verify that richer sparsity induces richer semantics, we here analyze pretrained models.
+In GPT-2 (Fig.
+4), orthant-based clustering recovers categories such as medical terms, emotional tone, entertainment, and political figures.
+In BERT (Fig.
+14 in App.), we find clusters reflecting numerical patterns (e.g., 3-digit numbers), action verbs, administrative phrases, and morphological forms.
+Applying the same method to Qwen-7B yields UI phrases, programming keywords, numeric tokens, functional Chinese expressions, and so on.
+Qwen requires combining more concepts (7–11 vs.
+5–7 for GPT-2) to isolate coherent groups (Table 1 in App.), suggesting more distributed semantic encoding, potentially reflecting its multilingual training.
+Consistent structure across models.
+Our analysis predicts that, as long as an architecture is sufficiently expressive for the UFM assumption to hold, the geometry of learned semantics should depend on data patterns rather than specific architectural details.
+We verify this by training two transformers with different depths (6 vs.
+12 layers) on the same Simplified TinyStories dataset.
+After training, we extract the decoder matrix $\bm{W}$} from each model and compute its top three left singular vectors.
+Applying orthant-based clustering, both models yield nearly identical semantic groupings (Fig.
+5).
 5 Related Work
-We bring together, in the context of NTP-trained language models, the following two classical ideas: (1) semantic patterns emerge from linear relationships between vector representations of textual inputs, and (2) these learned representations form geometries tied to language statistics via count matrices.
+We bring together, in the context of NTP-trained language models, two classical ideas: (1) semantics emerge from linear relationships between vector representations of textual inputs, and (2) learned-representations geometries are tied to language statistics via count matrices.
+App.
+B for details.
+Extracting semantics from distributed representations was popularized through word embedding research Mikolov et al.
+[2013a], with approaches including semantic axes An et al.
+[2018], Fast et al.
+[2016] and word analogies Mikolov et al.
+[2013b], Levy et al.
+[2015].
+In LLMs, work has focused on dictionary learning Turner et al.
+[2023], Li et al.
+[2024], geometric clustering Coenen et al.
+[2019], Hewitt and Liang [2019], and linear probing Hewitt and Manning [2019], Marks and Tegmark [2023].
+Unlike post-hoc analyses, we analyze, from an optimization perspective, how semantics emerge as a consequence of training.
+While Park et al.
+[2023, 2024] have explored similar questions, here we directly analyze NTP optimization, rather than relying on a generative model for concept generation and its link to words/contexts.
+Classical work has shown that applying factorization methods to co-occurrence matrices conveys semantic information Landauer [1997], Levy et al.
+[2015], with various hand-engineered choices for both the matrix representation (e.g., word-document, word-word, PMI, PPMI) and factorization technique (SVD, NNMF Lee and Seung [1999], Ding et al.
+[2006], sparse dictionary learning Murphy et al.
+[2012], Faruqui et al.
+[2015]).
+We demonstrate, for the first-time, that gradient-based optimization itself determines both the matrix representation ($\widetilde{\bm{S}}$) and factorization method (SVD) used to encode semantics.
+Our work connects NTP-UFM to the closed-form framework of Saxe et al.
+[2013, 2019] for linear neural networks.
+While their work focuses on general cognitive development with orthogonal inputs, we provide a new practical instantiation for language modeling optimization and validate theory findings on real data and architectures.
+6 Conclusion
+As LLM capabilities continue to advance, understanding their inner workings is crucial.
+We study how language-modeling optimization guides the geometry of word and context representations to organize around latent semantic information.
+Without explicitly computing it, the model encodes such information in the SVD factors of an implicit data-sparsity matrix.
+As an early contribution in this direction, our analysis motivates several future directions, some of which we outline in Sec.
+7.
+7 Limitations
+(1) While the assumption $d\geq V$ allows already for rich geometric structures, most practical models use $d<V$.
+We hypothesize that during NTP training, models learn to represent the $d$ most significant concepts, corresponding to the largest singular values of $\widetilde{\bm{S}}$.
+While NTP-UFM offers preliminary support of this hypothesis, a rigorous theoretical and experimental analysis requires separate future investigation.
+Encouragingly, since semantic categories arise from concept combinations (Sec.
+3.2), even a reduced set of core concepts could enable rich semantic representations.
+(2) Our analysis assumes sufficient expressivity and training time, which may become increasingly relevant with improved compute or in data-limited regimes, but relaxing them or experimentally validating their applicability in the spirit of our GPT-2/Qwen experiments is important.
+(3) Existing theoretical work on transformer optimization dynamics is limited to shallow models (e.g.
+one-layer) and overly strict data-generation assumptions that limit capturing meaningful semantic information.
+On the other hand, the abstraction of unconstrained features offers mathematical tractability, already useful insights, and has the attribute of leading to predictions insensitive to the specifics of the architecture (Sec.
+4.2).
+Future work can explore merging the two paradigms of analysis.
+(4) While our approach differs from concurrent efforts to explain semantic emergence in LLMs via probing or sparse-dictionary learning, future work could explore connections and potential applications to model improvement techniques.
+8 Acknowledgments
+This work was funded by the NSERC Discovery Grant No.
+2021-03677, the Alliance Grant ALLRP 581098-22, and an Alliance Mission Grant.
+The authors also acknowledge use of the Sockeye cluster by UBC Advanced Research Computing.
+Appendix A Paper Organization
+The main paper is organized as follows: Sec.
+2 introduces background on our analytical model and the geometry of embeddings.
+Our main analytical contributions and minimal synthetic experiments are presented in Sec.
+3, with subsections 3.13.3 each detailing one of the three contributions (C1-C3).
+Sec.
+4 provides qualitative studies on real-world data and pretrained models that validate our findings.
+Sec.
+5 contextualizes our work within the broader literature, with comparisons to closely related work integrated throughout the paper.
+We conclude in Sec.
+6 and discuss limitations and future directions in Sec.
+7.
+The appendix is organized as follows: Sec.
+B provides an extended discussion of related work (a condensed version appears in the main paper).
+Sec.
+C includes details on the formalization of concept representations discussed in Sec.
+3.1 and the proof of Theorem 1 from Sec.
+3.3.
+Additional experimental results deferred due to space constraints appear in Sec.
+D.
+Appendix B Extended Discussion on Related Work
+This section provides an extended discussion of the related work summarized in Sec. 5.
 Semantics from Distributed Representations.
@@
 [2011], Hu and Liu [2004], and (ii) investigating semantic patterns through “word analogies,” which reveal that semantic relationships manifest as linear directions in embedding space Mikolov et al.
-[2013c], Levy et al.
+[2013b], Levy et al.
 [2015], Drozd et al.
@@
 [2024].
-However, recent focus has shifted to techniques for distilling interpretable semantics through dictionary learning or sparse autoencoders Cunningham et al.
-[2023], Bricken et al.
-[2023], further showing these can be used for model steering Turner et al.
+However, recent focus has shifted to dictionary learning techniques for distilling interpretable semantics, e.g., Turner et al.
 [2023], Li et al.
-[2024] (techniques that in fact trace back to early work with static word embeddings, e.g., Murphy et al.
-[2012], Faruqui et al.
-[2015], Arora et al.
-[2018], Zhang et al.
-[2019]).
+[2024].
 Alternative approaches include geometric studies showing how linguistic features cluster in embedding spaces Coenen et al.
@@
 Simultaneously, supervised methods such as linear probing have been employed to extract semantic and syntactic information from modern LLMs Alain and Bengio [2017], Hewitt and Manning [2019], Marks and Tegmark [2023], relying on the same principle of linear representation found in word analogies.
-Unlike these aforementioned works that empirically analyze trained models post-hoc, we focus on formalizing the mechanisms by which semantic structure emerges during NTP training.
-Our goal is not to introduce a state-of-the-art method for extracting semantics, but rather to understand, from an optimization viewpoint, how semantics emerge as a consequence of next-token prediction.
+Unlike these works that empirically analyze trained models post-hoc, our goal is not to introduce a state-of-the-art method for extracting semantics, but rather to understand, from an optimization viewpoint, how semantics emerge as a consequence of next-token prediction.
 While Park et al.
-[2023, 2024] have explored similar questions, here, rather than relying on a model for concept generation and its link to words/contexts, we directly analyze NTP.
-Geometry from Co-occurrence Statistics.
-Vector space representations emerge from two approaches: (i) traditional count-based models using co-occurrence matrices, and (ii) prediction-based models that optimize token prediction objectives like Skip-gram or NTP.
+[2023, 2024] have explored similar questions, here we directly analyze NTP optimization, rather than relying on a generative model for concept generation and its link to words/contexts.
+Geometry from Co-occurrence Statistics and Matrix Factorization.
+The idea that factorizations of co-occurrence matrices convey latent semantic information dates back to latent semantic analysis of word-document co-occurrence matrices Landauer [1997], Deerwester et al.
+[1990] and later to pointwise mutual information matrices of word-word co-occurrences Levy and Goldberg [2014], Levy et al.
+[2015].
+Beyond SVD, alternative factorization techniques such as non-negative matrix factorization Lee and Seung [1999], Ding et al.
+[2006] and sparse dictionary learning Murphy et al.
+[2012], Faruqui et al.
+[2015] have been applied to various transformations of co-occurrence data, including those addressing PMI sparsity (e.g., PPMI) or using probability smoothing Levy et al.
+[2015].
+More broadly, vector space representations emerge from two approaches: (i) traditional count-based models using co-occurrence matrices, and (ii) prediction-based models that optimize token prediction objectives like Skip-gram or NTP.
 Levy and Goldberg [2014], Pennington et al.
@@
 [2024] have extended this insight to causal NTP models with a key distinction: the co-occurrence matrix between contexts and words is inherently sparse, unlike the dense word-word co-occurrences in prior work.
-This distinction makes relevant recent advances in implicit regularization Ji and Telgarsky [2019], Ji et al.
+This distinction makes relevant recent advances in implicit regularization Soudry et al.
 [2018] and neural collapse geometries Papyan et al.
 [2020b], Mixon et al.
-[2020], Fang et al.
-[2021], through which Zhao et al.
-[2024] analyze how word and context embeddings converge via an unconstrained features model, relating them to singular factors of the centered co-occurrence probability matrix’s support set (see Eq.
-(2)).
-We show that these singular factors encode rich semantic structure.
-While we argue that the natural language’s inherent multilabel nature is responsible for the emergence of rich semantics, making full circle to the neural-collapse literature, we show that the semantic interpretation applies to as simple a setting as one-hot imbalanced classification.
-Moreover, by connecting to the closed-form analysis of learning dynamics of linear neural networks by Saxe et al.
-[2013, 2019], Gidel et al.
-[2019], we show how singular values determine the learning order during training.
-More details on related works deferred to App. E. See also Remark 1.
-3 Background
-Notations.
-For any integer $k$, $[k]:=\{1,\ldots,k\}$.
-Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$ respectively.
-For matrix $\bm{A}$, $\bm{A}[i,j]$ is its $(i,j)$-th entry, and for vector $\bm{a}$, $\bm{a}[i]$ is its $i$-th entry.
-$\mathds{1}$ is the all-ones vector and $\otimes$ is Kronecker product.
-3.1 NTP Objective as Sparse Soft-Label Classification
-[…]
-$\mathcal{L}({\bm{W}},\bm{\theta})=\sum\nolimits_{j\in[m]}\hat{\pi}_{j}\cdot\ell\left({\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j});\hat{\bm{p}}_{j}\right)\,,$ (1)
-where the loss $\ell$ penalizes deviations between the model’s logits ${\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j})$ for context $j$ and its corresponding soft-label vector $\hat{\bm{p}}_{j}$.
-Unless otherwise specified, we use the standard cross-entropy loss we set the loss $\ell$ to be the standard CE loss, i.e., $\operatorname{CE}\left(\mathbb{S}({\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j}))\,||\,\bm{p}_{j}\right)$, where $\mathbb{S}()$ denotes the softmax map of logits to the probability simplex.
-For later use, define the support matrix $\bm{S}\in\{0,1\}^{V\times m}$ of the conditional probability matrix ${\bm{P}}=[\hat{\bm{p}}_{1},\ldots,\hat{\bm{p}}_{m}]\in\mathbb{R}^{V\times m}$.
-Formally, $\bm{S}[z,j]=1\,\Leftrightarrow\,{\bm{P}}[z,j]>0$.
-For context $j$, we refer to tokens $z$ for which $\bm{S}[z,j]=1$ as in-support tokens, and refer to all others as off-support.
-Central to our analysis is the centered support matrix
-$\widetilde{\bm{S}}:=(\mathds{I}_{V}-\frac{1}{V}\mathds{1}_{V}\mathds{1}_{V}^{\top})\bm{S}\,.$ (2)
-For convenience, we refer to $\widetilde{\bm{S}}$ as the data-sparsity matrix, though note that unlike the support matrix $\bm{S}$, its entries are not binary due to the centering operation:
-$\widetilde{\bm{S}}[z,j]=\begin{cases}1-\nicefrac{{|{\mathcal{S}}_{j}|}}{{V}}&\text{if $\bm{S}[z,j]=1$}\,,\\ -\nicefrac{{|{\mathcal{S}}_{j}|}}{{V}}&\text{if $\bm{S}[z,j]=0$}\,.\end{cases}\,$
-3.2 Geometry of Words and Contexts
-Following Yang et al.
-[2017], we assume the model is sufficiently expressive to allow optimizing context embeddings in (1) freely, instead of abiding by their architecture-specific parameterization.
-Concretely, this modeling abstraction yields a simplified training objective
-$\min\nolimits_{{\bm{W}},{\bm{H}}}~{}~{}\mathcal{L}({\bm{W}}{\bm{H}})+\frac{{\lambda}}{2}\|{\bm{W}}\|^{2}+\frac{{\lambda}}{2}\|{\bm{H}}\|^{2}\,,$ (NTP-UFM)
-which jointly optimizes ${\bm{W}}\in\mathbb{R}^{V\times d}$ and ${\bm{H}}:=[\bm{h}_{1},\ldots,\bm{h}_{m}]\in\mathbb{R}^{d\times m}$, where $\bm{h}_{j}$ is the freely optimized embedding of context $\bar{\bm{x}}_{j}$.
-Since the minimization is unconstrained for both variables, we follow Mixon et al.
-[2022], Wojtowytsch [2021], Fang et al.
-[2021] in referring to this as the unconstrained features model (UFM) for NTP training.
-We refer to ${\bm{W}}$ and ${\bm{H}}$ as the word and context embedding matrices.
-Eq.
-(NTP-UFM) includes ridge-regularization with weight $\lambda>0$.
-Zhao et al.
-[2024] have analyzed the geometry of solutions to (NTP-UFM) when $\lambda\rightarrow 0$.
-This limit (referred to in the literature as the regularization-path) serves as a proxy for the limiting behavior of gradient descent (GD) training as the number of iterations approaches infinity Ji et al.
 [2020].
-For large embedding dimensions $d\geq V$,  [^2] See Sec.
-A.1 for a discussion on this assumption.
-as $\lambda\rightarrow 0$, Zhao et al.
-[2024] show the following:
-1.
-Logits Convergence: The logit matrix $\bm{L}={\bm{W}}{\bm{H}}$ decomposes into two orthogonal components: one that grows unboundedly during training and another that remains finite.
-After normalization (required to characterize the convergence of diverging logits), only the unbounded component becomes dominant, denoted as $\bm{L}^{\rm{mm}}$.
-Crucially, $\bm{L}^{\rm{mm}}$ depends only on the data support matrix $\bm{S}$ (see Zhao et al.
-[2024] for details).
-2.
-SVD factors of $\bm{L}^{\rm{mm}}$: Like logits, word and context embeddings grow unboundedly in magnitude, but converge in direction to ${\bm{W}}^{\rm{mm}}=\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ and ${\bm{H}}^{\rm{mm}}=\mathbf{R}^{\top}\bm{\Sigma}^{1/2}{\bm{V}}^{\top}$, respectively.
-Here, $\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ is the SVD of $\bm{L}^{\rm{mm}}$ and $\mathbf{R}$ is a partial orthogonal matrix.
-3.
-Data-sparsity matrix as proxy: The matrix $\widetilde{\bm{S}}$ (see Eq.
-(2)) is a good proxy for $\bm{L}^{\rm{mm}}$.
-[^3] This is formally proven in symmetric cases and supported by extensive empirical evidence.
-Put together, the singular factors of $\widetilde{\bm{S}}$ play a crucial role in determining the geometry of word and context embeddings learnt by NTP training.
-Denote this SVD as
-$\widetilde{\bm{S}}:=\bm{U}\bm{\Sigma}{\bm{V}}^{\top},\quad$ (3)
-where $\bm{U}\in\mathbb{R}^{V\times r}$, ${\bm{V}}\in\mathbb{R}^{m\times r}$ with $\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}\,,$ the singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered: $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}>0\,$, and $r\leq V-1$ denotes the rank.
-4 Geometry of Concepts
-4.1 Motivating Questions
-While prior work explicitly describes the geometry of word and context embeddings, it leaves open two key questions: (i) What is the latent information that drives the word/context embeddings to this geometry?
-And, (ii) does the resulting geometry form structures (e.g.
-in the form of clusters) that lend themselves to human-interpretable semantics?
-To illustrate this, Fig.
-2A shows a 3D projection of trained word and context embeddings on a simplified dataset for visualization.
-The conditional probability matrix ${\bm{P}}$ is shown on the right.
-The geometry reveals clear structure: animal-related words and contexts cluster on the left, plant-related ones on the right.
-This suggests the emergence of latent concept information—specifically, a positive/negative semantic distinction.
-From Sec.
-3.2, we know word embeddings take the form ${\bm{W}}=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$, where $\bm{U}$ encodes the row-space of $\widetilde{\bm{S}}$ (Eq.
-(2)).
-Since rows of $\widetilde{\bm{S}}$ correspond to words, and words that appear in similar contexts (e.g., “the organism that grows roots is…” vs.
-“the organism that has green leaves is…”) have similar patterns, $\bm{U}$ may naturally encode semantic relationships.
-Its columns can be interpreted as a “concept basis”, where a word’s projection reflects its association with these latent concepts.
-4.2 Principal Components as Concepts
-We term each dimension $k\in[r]$ resulting from the SVD of $\widetilde{\bm{S}}$ as a concept.
-Adopting terminology from Saxe et al.
-[2019], we call the columns $\mathbf{u}_{k}\in\mathbb{R}^{V}$ and $\mathbf{v}_{k}\in\mathbb{R}^{m}$ of matrices $\mathbf{U}$ and $\mathbf{V}$ the “word analyzer vectors” and “context analyzer vectors,” respectively.
-These vectors represent the alignment of tokens and contexts with each concept: for word $z$ and concept $k$, the sign of $\mathbf{u}_{k}[z]$ indicates association ($\mathbf{u}_{k}[z]>0$) or opposition ($\mathbf{u}_{k}[z]<0$) to the concept, while its magnitude quantifies the strength of this relationship.
-[^4] Note that each word analyzer vector $\bm{u}_{k},k\in[r]$ has both positive and negative entries (corresponding to association and opposition), because $\bm{u}_{k}$ is orthogonal to $\mathds{1}_{V}$ (which is in the nullspace of $\widetilde{\bm{S}}$).
-Words for which $|\bm{u}_{k}[z]|$ is small are neutral to this concept.
-The same interpretation applies to context components $\mathbf{v}_{k}[j]$, measuring each context’s alignment with the respective concept.
-Fig.
-1 verifies the SVD of $\widetilde{\bm{S}}$ captures the dataset’s semantic information for the simplified [^5] This design isolates semantic concept extraction by minimizing grammatical influences.
-These simplifications are for visualization clarity only—the principles apply to complex datasets with combined grammatical/semantic attributes and to full autoregressive training (see Secs.
-5 and A.2).
-dataset with syntax "The organism that [attribute] is [subject]" (see Fig.
-2).
-To represent concepts in the embedding space and thus directly characterize their relationships with word and context embeddings, we define the $d$-dimensional “word-concept representations” $\bm{u}^{\rm{d}}_{k}$ and “context-concept representations” ${\bm{v}}^{\rm{d}}_{k}$ for $k\in[r]$ as projections onto the respective spaces of words and contexts, that is (see also Sec.
-B for details), $\bm{u}^{\rm{d}}_{k}={\bm{W}}^{\top}\bm{u}_{k}$ and ${\bm{v}}^{\rm{d}}_{k}={\bm{H}}{\bm{v}}_{k}.$ This ensures that words or contexts more closely aligned with a specific concept have embeddings closer to the concept’s representation.
-Similar to classical semantic axes formed by averaging relevant words, these representations are weighted averages where weights ($\bm{u}_{k}$) reflect concept relevance.
-Fig.
-10 visualizes two most significant concept representations and their relationships with word/context embeddings.
-Remark 1.
-The idea that SVD factors of a co-occurrence-type matrix convey latent semantic information dates back at least to latent semantic analysis of word-document co-occurrence matrices Landauer [1997], Deerwester et al.
-[1990] and later to pointwise mutual information matrices of word-word co-occurrences Levy and Goldberg [2014], Levy et al.
+Relation to [Zhao et al., 2024].
+Our work extends Zhao et al.
+[2024] by addressing the three open questions Q1-Q3 discussed in our Introduction.
+In doing so, we provide crucial semantic interpretation to prior NC-based representation analyses (including those for imbalanced one-hot settings Fang et al.
+[2021], Thrampoulidis et al.
+[2022a]).
+Furthermore, by connecting our model to the closed-form analysis of linear neural network learning dynamics Saxe et al.
+[2013], we rigorously demonstrate how singular values govern the order of concept acquisition during training.mine the learning order during training.
+Relation to Saxe et al.
 [2019].
-Beyond SVD factorization, researchers have explored alternative factorization techniques, such as non-negative matrix factorization Lee and Seung [1999], Ding et al.
-[2006] and sparse dictionary learning Murphy et al.
-[2012], Faruqui et al.
-[2015], and have applied them to various transformations of co-occurrence data, including modifications that address the sparsity of PMI matrices by eliminating negative entries (e.g., PPMI) or applying probability smoothing techniques Levy et al.
-[2015].
-Our key conceptual contribution is that we let the training dynamics NTP optimization determine both the matrix representation of the data and the factorization method used to encode latent semantics: the model naturally utilizes the centered support matrix $\widetilde{\bm{S}}$ and its SVD factorization.
-We validate through structured examples in this section and through experiments in Sec.
-5, that the model’s emergent representational choices successfully encode linguistic components without explicit engineering of the representation space.
-4.3 Orthant-based Clustering in Singular Spaces
-Concepts extracted from $\widetilde{\bm{S}}$, represent significant components from data.
-However, as noted, for example, by Chersoni et al.
-[2021], Piantadosi et al.
-[2024], not all concepts individually correspond to linguistically interpretable factors.
-Instead, we posit that human-interpretable semantics can be identified by specific combinations of several concepts.
-Here, we formalize this hypothesis by relating it to an orthant-based clustering of word/context representations.
-Recall from Sec.
-4.2 that the $k$-th word analyzer vector $\bm{u}_{k}$ categorizes words $z$ based on $\texttt{sign}(\bm{u}_{k}[z])$: words with large $|u_{k}[z]|$ exhibit strong correlation with concept $k$, with $\texttt{sign}(\bm{u}_{k}[z])$ indicating whether this correlation is positive/negative.
-We extend this categorization beyond single dimensions to multiple concept dimensions simultaneously.
-Let $\mathcal{K}=\{k_{1},\ldots,k_{p}\}$ denote a set of $p\leq r$ selected dimensions $k_{i}\in[r],i\in[p]$.
-Given $\mathcal{K}$, we define $2^{p}$ possible signature configurations $C=C(\mathcal{K})=[c_{k_{1}},\ldots,c_{k_{p}}]\in\{\pm 1\}^{p}$, where $c_{k_{i}}\in\{\pm 1\}$ indicates whether dimension $k_{i}$ contributes positively or negatively to the semantic category.
-Each signature configuration represents a potential semantic category, characterized by its member words and their degrees of association with the category.
-Definition 1.
-For $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ with sign configurations $c_{k_{i}}\in\{\pm 1\}$ for dimensions $k_{i}\in[r]$:
-  • A word $z$ is a member of $C$ (denoted $z\in C$) if and only if $\texttt{sign}(\bm{u}_{k_{i}}[z])=c_{k_{i}}$ for all $k_{i},i\in[p]$.
-  • The typicality of a member $z\in C$ is defined as: $\text{Typicality}(z;C)=\sum\nolimits_{i\in[p]}\big{|}\bm{u}_{k_{i}}[z]\big{|}\,.$
-Analogous definitions hold for contexts.
-Geometrically, each configuration corresponds to a $p$-dimensional orthant (or orthant sheet) in the $d$-dimensional embedding space.
-Membership indicates which word/context embeddings lie within this orthant, effectively creating an orthant-based clustering in the subspace spanned by the selected concept dimensions $\mathcal{K}$.
-The typicality of a member measures the L1-norm (other choices are also possible) of its representation’s projection (the $z$-th row of $\bm{U}$ for words, $j$-th row of ${\bm{V}}$ for contexts) onto the selected concept dimensions.
-Fig.
-2B and Sec.
-5 demonstrate that this orthant-based clustering reveals interpretable semantic categories, while also inducing a hierarchical structure, where the number of concept dimensions determines the semantic granularity: broad linguistic categories (e.g., verbs) require fewer dimensions, while more specific categories (e.g., past-tense verbs) need more dimensions for precise characterization.
-See also Sec.
-C and Fig.
-11 in the appendix.
-4.4 Rate of Learning
-We show that the hierarchical structure of semantic concepts discussed in Sec.
-4.3 is reflected in their emergence during training: broad semantic distinctions are learned before finer-grained features.
-Specifically, we track the singular values of the logit matrix $\bm{L}_{t}={\bm{W}}_{t}{\bm{H}}_{t}$ at training iteration $t$ when minimizing the unconstrained (${\lambda}\rightarrow 0$) objective (NTP-UFM).
-On the one hand, at the end of training ($t\rightarrow\infty$), singular values exhibiting the largest divergence correspond to the largest singular values of $\widetilde{\bm{S}}$ Zhao et al.
-[2024], which represent the most important concepts.
-On the other hand, at initialization, all singular values of $\bm{L}_{t}$ are small (following typical small-scale initialization).
-As training progresses, Fig.
-18 shows that the singular values increase at different rates: those that ultimately achieve the greatest magnitudes begin diverging earlier.
-Fig.
-3 provides empirical evidence of this hierarchical concept acquisition.
-The confusion matrices from a 2-layer transformer trained on natural language data reveal that the model first masters broad categorical distinctions (e.g., "Animals" vs.
-"Plants") before progressing to more specific semantic features.
-This progression hints at a coarse-to-fine learning order, which we formalize in the following theoretical analysis.
-True ClassEpoch 0Refer to captionEpoch 20Refer to captionEpoch 40Refer to captionEpoch 200Refer to captionPredicted Class
-Figure 3: Confusion matrices showing the progression of concept learning in a 2-layer transformer (embedding dimension $d=128$) across training epochs (0, 20, 40, 200).
-Each matrix cell shows the number of predictions for each true class.
-Classes correspond to support sets defined in Fig.
-2C, with semantic labels (e.g., "Animals", "Plants") given in brackets where applicable.
-For example, the support set dog, wolf, cat, tiger, sparrow, eagle, trout, shark represents the "Animals" class (first row/column).
-The matrices illustrate a temporal learning hierarchy: broad categories like "Animals" and "Plants" emerge by epoch 20, while finer-grained concepts (e.g., "Reproduces with pollen", "Lives in water") develop later, between epochs 40–200.
-This reflects that high-variance (large singular value) concepts—often aligned with broader semantics—are learned earlier during training.
-To gain theoretical insights into this behavior, we analyze (NTP-UFM) with square loss.
-[^6] This choice is motivated by the technical challenges in finite-time characterization of GD dynamics for CE loss due to coupling between singular directions.
-The simplified, square loss, could be seen as a first-order approximation of CE loss around zero initialization or in the high-temperature limit.
-Also, while CE loss is predominant in practice, we note that square-loss-type objectives have demonstrated practical success, as in the GloVe model Pennington et al.
-[2014], which achieves performance comparable to word2Vec Mikolov et al.
-[2013d].
-We hope that future work can prove the analogous phenomena we observe empirically for CE loss.
-Specifically, we show in Sec.
-D in the appendix that the NTP-UFM setting is equivalent to the setting studied extensively by Saxe et al.
-[2013, 2019], Gidel et al.
-[2019]; thus, with sufficiently small initialization and learning rate, the $k$-th singular value of $\bm{L}_{t}$ undergoes a sharp phase transition from (almost) 00 to its limit $\sigma_{k}$ at time $T_{k}\propto 1/\sigma_{k}$ (see Thm.
-1).
-In other words, more significant concepts (corresponding to larger singular values) are learned first, i.e.
-$T_{1}\leq T_{2}\leq\cdots\leq T_{r}$.
-We further validate this behavior in a controlled imbalanced setting with one-hot labels.
-Two majority classes each appear across many contexts, while two minority classes appear in only one.
-As shown in Figure 9, the model initially predicts all inputs as the dominant majority class (Step 0), then begins separating the majority classes (Step 13), followed by gradual recognition of the minority classes (Step 30), and resolves confusion between them later in training (Step 55).
-This progression reflects the emergence of concepts in order of their singular values, consistent with theoretical analysis.
-5 Experiments
-5.1 Setup
-Datasets.
-We conduct experiments on both synthetic data and pretrained models.
-For our synthetic experiments, we constructed two datasets by subsampling from two distinct corpora: TinyStories Eldan and Li [2023] and WikiText-2 Merity et al.
-[2017].
-For both datasets, we fix the vocabulary size to $V=1,000$ and extract $m=10,000$ contexts, creating the "Simplified TinyStories/WikiText" datasets.
-We extracted the contexts by sampling frequent contexts of lengths 2-6 from the entire datasets.
-This variation in context-length emulates the autoregressive nature of training with sliding window, which yields rich data-sparsity patterns that in turn result in richer semantic information (compared to the last-token visuals in Sec.
-4).
-For our experiments on pretrained transformers, we chose GPT-2 Radford et al.
-[2019] and extracted $50,000$ text sequences, each 11 tokens in length from GPT-2’s training corpus, the OpenWebText dataset.
-By segmenting these sequences into contexts of varying lengths (1-10 tokens), we generated $500,000$ contexts, which we designate as the "simplified OpenWebText" dataset.
-Concepts and semantics.
-Following Sec.
-4, for the simplified datasets where it is computationally feasible, we construct the $V\times m$ data-sparsity matrix $\widetilde{\bm{S}}$ and compute its SVD.
-The left/right singular vectors serve as our word/context concepts.
-As per Sec.
-4.3, we also consider signature-based concept combinations, which we examine for semantic interpretability as discussed in Visuals below.
-For completeness, we also experiment with $k$-means spectral clustering as an alternative way to derive interpretable semantics from top important concepts.
-For the OpenWebText dataset, instead of performing SVD on $\widetilde{\bm{S}}$ (which is computationally expensive to construct and factorize), we work directly with the embeddings of the pretrained model.
-For word concepts, we extract ${\bm{W}}$ from GPT-2’s decoder and take the matrix $\bm{U}$ of word concepts to be its left singular matrix.
-For context concepts, we form a $d\times m$ matrix ${\bm{H}}$ by concatenating the embeddings (last-MLP layer representations before the decoder) of GPT-2 for each context in our subset.
-This matrix corresponds to only a portion of the “true” embedding matrix for the entire dataset.
-To ensure the coupling between ${\bm{W}}$ and ${\bm{H}}$ that is inherent to the SVD of $\widetilde{\bm{S}}$ (which we don’t have direct access to), we compute the matrix ${\bm{V}}$ of context concepts as ${\bm{V}}={\bm{H}}^{\top}\mathbf{R}\bm{\Sigma}^{-1/2}$, where $\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ denotes the SVD of ${\bm{W}}$ that we already computed.
-Visuals.
-To interpret the semantic information, we visualize the top $40$ words/contexts associated with each concept.
-“Top” here is measured with respect to the typicality score in Defn.
-1.
-For visualization, we use word clouds where size is proportional to its typicality score.
-We often use the pound sign (#) to indicate token start.
-5.2 Findings
-Individual concepts are not interpretable.
-We have seen in Sec.
-4 that not all concepts individually correspond to linguistically interpretable factors.
-For completeness, we verify this conclusion for the Simplified TinyStories/WikiText datasets in Fig.
-14 in the appendix.
-This observation is well-reported in the literature, e.g., Chersoni et al.
-[2021], and is also related to the recently popularized superposition hypothesis Elhage et al.
-[2022].
-Combining concepts with signature-configurations reveals semantics.
-We demonstrate that combining non-interpretable concepts can reveal human-interpretable semantic information.
-Here, we show this by implementing the idea of orthant-based clustering of Sec.
-4.3.
-Specifically, we combine $p=5$ to $7$ concepts in successive order starting from the most important one, i.e.
-$\mathcal{K}=\{1,\ldots,p\}$.
-Fig.
-4 shows human-interpretable semantics identified by orthant-based clustering in Simplified TinyStories and WikiText-2 datasets.
-In Fig.
-11, we show how orthant-based clustering progressively reveals hierarchical semantic structure as we combine more concept dimensions, illustrated through the example of verb forms organized by grammatical distinctions.
-Starting with a single dimension ($\mathcal{K}=\{1\}$), the categories remain broad and difficult to interpret.
-However, as we incorporate additional dimensions, increasingly refined semantic categories emerge, ultimately distinguishing between different verb forms such as past/present tense, and present continuous forms.
-Past-tense verb
-Present-tense verb
-Preposition
-Proper names
-Month
-Number
-Unit
-Preposition/Conjunction
-Figure 4: Semantic clusters based on signature configurations of Sec.
-4.3.
-Top: Simplified TinyStories dataset showing Past-tense verbs, Present-tense verbs, Prepositions, and Proper names.
-Bottom: Simplified Wikitext dataset showing Months, Numbers, Units, and Prepositions/Conjunctions.
-Each configuration is shown in its title.
-Past-tense verbs Refer to caption
-Proper names Refer to caption
-Numbers Refer to caption
-Conjunctions Refer to caption
-Figure 5: $k$-means cluster with top-$p=\log_{2}(k)$ dimensions of word-analyzer vectors.
-Left: 2 of 32-means clusters with $\mathcal{K}=\{1,\ldots,5\}$ on Simplified TinyStories representing Past-tense verbs and Proper names; Right: 2 of 64-means cluster with $\mathcal{K}=\{1,\ldots,6\}$ on simplified WikiText, representing Numbers and Conjunctions.
-$k$-means spectral clustering.
-As a complementary approach to our orthant-based clustering, we also investigate k-means spectral clustering.
-Specifically, we perform k-means clustering according to pairwise distances of analyzer vectors with their top $p=\log_{2}(k)$ dimensions (corresponding to top-$p$ most important concepts).
-Vanilla $k$-means spectral clustering selects $k$ principal dimensions.
-The modification here is inspired by the orthant-based clustering for which $p$ concepts correspond to $2^{p}$ orthants.
-Fig.
-5 shows semantic meaning in a subset of the recovered clusters with $k=32$ and $k=64$ for Simplified TinyStories and WikiText-2, respectively; see Sec.
-F.5 for full results.
-See also discussion in Sec.
-A.3.
-Relation of words and context semantics.
-Fig.
-2A demonstrates a relationship between context embeddings and word embeddings that share semantic features: contexts and words representing the same semantic concept (e.g., plant or animal) cluster proximally in the embedding space.
-This pattern extends to the richer datasets we study here.
-To verify this, we perform orthant-based clustering of contexts and words based on the same signature configuration of a subset of concept dimensions.
-Fig.
-6 shows this for Simplified TinyStories: contexts containing modal constructions like "want/like/love/have to" correspond strongly with word clusters containing infinitive verbs.
-Similarly, contexts with intensifier patterns "was so/very" align with clusters of descriptive adjectives.
-This relation reflects a natural dual semantic interpretation of contexts: they can be understood either through the direct semantic meaning of their constituent words or through the semantic properties of words that typically follow them.
-The latter interpretation aligns naturally with the predictive nature of NTP, where contexts acquire meaning partly from what they predict.
-  Refer to caption Refer to caption
-Figure 6: Context-word cluster pairs sharing identical sign patterns across top 5 concepts.
-Left and right pair illustrates the “verb+to infinitive” and “was so/very + adj” structures, confirming semantic alignment between contexts and their in-support words.
-Semantics in pretrained models.
-We find interpretable semantics from simplified datasets are primarily grammatical or syntactic due to limited sample size and context simplicity.
-While these features make analytical computations feasible (e.g., computing/factorizing $\widetilde{\bm{S}}$), they compromise the richness of sparsity patterns and, in turn, the emergence of semantics.
-To verify that richer sparsity patterns yield richer semantics, we additionally experiment with pre-trained GPT-2.
-See Fig.
-7 and Sec.
-F.2 in appendix for additional results.
-Positive
-Negative
-Chemical
-Medical
-Coorperation
-Digital entertainment
-Cultural
-Societal
-Figure 7: Semantics identified by orthant-based clustering on GPT2’s word embeddings.
-6 Conclusion and Limitations
-While continuing to advance LLMs’ capabilities, understanding their inner workings is crucial.
-We study how NTP optimization guides the geometry of word and context representations to organize around latent semantic information.
-Without explicitly computing it, the model encodes such information in the SVD factors of an implicit data-sparsity matrix.
-As an early contribution in this direction, our analysis motivates several future directions: (1) While the assumption $d\geq V$ allows for rich geometric structures, a formal investigation is needed for the $d<V$ case prevalent in practice (see Sec.
-A.1).
-(2) Our analysis assumes sufficient expressivity and training time, which may become increasingly relevant with improved compute or in data-limited regimes, but relaxing them or experimentally validating their applicability in the spirit of our GPT-2 experiments (e.g., in the spirit of Wu and Papyan [2024]) is important.
-(3) Understanding how transformer architectures specifically fit into our analytical framework remains an open challenge.
-(4) Some conclusions rely on square loss and our work highlights a significant gap in understanding CE loss dynamics, even in simple bilinear models.
-(5) While our approach differs from concurrent efforts to explain semantic emergence in LLMs via probing or sparse-dictionary learning, future work could explore connections and potential applications to model improvement techniques.
-7 Acknowledgments
-This work was funded by the NSERC Discovery Grant No.
-2021-03677, the Alliance Grant ALLRP 581098-22, and an Alliance Mission Grant.
-The authors also acknowledge use of the Sockeye cluster by UBC Advanced Research Computing.
-Appendix A Additional Discussions
-A.1 On the Assumption $d\geq V$
-Our analysis has assumed the embedding dimension $d$ is at least as large as the rank of the data-sparsity matrix (guaranteed when $d\geq V$), allowing word and context embeddings to form geometric structures representing all concepts encoded in the matrix’s singular factors Zhao et al.
-[2024].
-As noted in Zhao et al.
-[2024], while this assumption differs from current practice in LLMs, the geometry of embeddings remains rich in this setting.
-Importantly, this assumption is less restrictive than requiring $d>C$ in one-hot classification with $C$ classes.
-There, due to collapse of embeddings from the same class, $d>C$ effectively requires the dimension to exceed the number of training examples.
-In contrast, for NTP, the number of contexts $m$ can be (and typically is) much larger than $d$, allowing for rich geometric arrangements of context embeddings in the lower-dimensional space and non-trivial emergence of concepts, which is confirmed throughout our case studies.
-That said, the current trend of modern language models typically employing $d<V$, raises the question: Which concepts are prioritized when the model cannot capture all of them?
-We hypothesize that during NTP training, models learn to represent the $d$ most significant concepts, corresponding to the largest singular values of the data-sparsity matrix.
-While this requires a separate study that is beyond our current scope, our preliminary investigation using NTP-UFM with square-loss on a synthetic dataset supports this hypothesis.
-Fig.
-19 shows the learned logit matrix’s singular values converging to the $d$ largest singular values of the data-sparsity matrix during training.
-Following Sec.
-4.4, we used square loss rather than CE to ensure bounded singular values.
-While a rigorous analysis for $d<V$ remains future work, since meaningful semantic categories emerge from combinations of concepts, even a reduced set of core concepts could enable rich semantic representations through the orthant clustering mechanism of Sec.
-4.3.
-A.2 Role of Autoregression
-We have shown that concepts emerge during NTP training as principal directions of $\widetilde{\bm{S}}$, a centered version of the support matrix $\bm{S}$.
-Each column $j$ of $\bm{S}$ corresponds to a context and can be viewed as a binary (multi-)label vector $\bm{s}_{j}$, where entries of $1$ indicate tokens that appear as next-tokens for that context in the training data.
-While concepts are determined by the principal directions of this label matrix, this might seem limiting: contexts can relate not only through shared next-tokens but also through their intrinsic structure (e.g., overlapping constituent tokens).
-The autoregressive nature of training naturally captures these structural relationships as follows.
-As the model processes labels for progressively longer contexts, each context $(z_{1},\ldots,z_{t})$ contributes to concept formation both through its distribution over next-tokens $z_{t+1}$ and through the labels of its shorter subsequences $(z_{1},\ldots,z_{t-1})$, $(z_{1},\ldots,z_{t-2})$, etc.—each representing distinct columns in $\bm{S}$.
-This overlapping window structure produces fine-grained label information, creating rich sparsity patterns in $\bm{S}$ (and consequently in $\widetilde{\bm{S}}$) that yield nontrivial concepts with varying significance, as reflected in word and context embeddings.
-A.3 Relationship to Spectral Clustering
-The orthant-based clustering approach introduced in this work shares foundational similarities with spectral clustering while providing distinct advantages for linguistic interpretation.
-Both methods operate in reduced-dimensional spaces and leverage embedding geometry to identify meaningful clusters.
-First, while both approaches leverage singular factors, the orthant-based method provides a more direct link (in view of concept definition in Sec.
-4) between NTP and the emergence of semantic structure by explicitly interpreting the sign patterns of SVD dimensions as semantic indicators.
-An interesting property of our method is its selective nature: while $p$ dimensions generate $2^{p}$ potential orthants, only a subset contains semantically coherent word groupings.
-This selectivity reflects linguistic reality, where not all mathematical combinations yield meaningful semantic categories.
-In contrast, spectral clustering typically partitions the space completely, which may not always align with natural semantic boundaries in language.
-Finally, the orthant-based method offers semantic interpretation of membership and typicality, and dimensionality control through parameter $p$.
-This aligns with language model training dynamics, as the most important singular value components are learned first during training, making the resulting clusters reflective of the model’s knowledge acquisition process.
-To this respect, note that unlike orthant-based clustering producing a potential maximum of $2^{p}$ clusters when selecting $p$ top concepts, typical clustering algorithms such as vanilla $k$-means on principal components would result in $p$-clusters with the same number of selected concepts.
-Thus, when we evaluate spectral methods in our experiments, we choose $k$ on the order of $2^{p}$ for the value of $p$ that we find orthant-based clustering reveals semantically informative orthants.
-A.4 Neural Collapse Meets Semantics
-The emergence of concepts with rich semantic meanings stems from the interplay of labels (encoded in $\bm{S}$ or its centered version $\widetilde{\bm{S}}$) across different contexts.
-Consider the minimal label richness case: each context has exactly one next-token, with contexts distributed equally across the vocabulary.
-Here, $\bm{S}$ can be rearranged as $\mathds{I}_{V}\otimes\mathds{1}_{m/V}^{\top}$, yielding trivial concepts where all singular directions contribute equally.
-This setting parallels standard balanced one-hot classification studied in neural collapse literature, where last-layer embeddings and weights form highly symmetric aligned structures Papyan et al.
-[2020a], reflecting the symmetric nature of $\bm{S}$ where labels induce no interesting conceptual structure.
-However, such balanced settings never occur in autoregressive NTP, which as discussed in Sec.
-A.2, produces rich label formations.
-In these richer settings, $\bm{S}$ induces a geometry of concepts that provides semantic interpretation to the embedding geometry studied in the neural collapse literature.
-To illustrate this new perspective, we demonstrate that meaningful concepts emerge even in one-hot classification—the traditional focus of neural collapse literature—given minimal deviation from perfect balance.
-Specifically, consider STEP imbalances with ratio $R$: $V/2$ majority classes each have $R>1$ times more samples than minority classes.
-Here, Thrampoulidis et al. [2022] shows the learned logit matrix converges to $\widetilde{\bm{S}}$, whose singular values exhibit a three-tier structure:
-$\sigma_{1}=\ldots=\sigma_{V/2-1}=\sqrt{R}>\sigma_{V/2}=\sqrt{(R+1)/2}>\sigma_{V/2+1}=\ldots=\sigma_{V}=1.$ The left singular vectors matrix $\bm{U}$ takes a sparse block form:
-$\bm{U}=\left[\begin{array}[]{ccc}\mathbb{F}&-\sqrt{\frac{1}{V}}\mathbf{1}&\mathbf{0}\\ \mathbf{0}&\sqrt{\frac{1}{V}}\mathbf{1}&\mathbb{F}\end{array}\right]\in\mathbb{R}^{V\times(V-1)}\,,$
-where $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ is an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$.
-[^7] For concreteness, $\mathbb{F}$ can be constructed using the discrete cosine transform matrix, excluding the constant column: $\mathbb{F}[i,j]=\sqrt{\frac{4}{V}}\cdot\cos\left(\frac{\pi(2i-1)j}{V}\right)$ for $i\in[V/2]$, $j\in[V/2-1]$ The structure of $\bm{U}$ reveals three distinct types of concepts, corresponding to the three tiers of singular values: (1) First $V/2-1$ columns (non-zero only for majority classes) represent distinctions among majority classes; (2) Middle column (singular value $\sqrt{(R+1)/2}$) has opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy; (3) Last $V/2-1$ columns (non-zero only for minority classes) capture distinctions among minority classes.
-This hierarchical structure shows the network learns concepts in order: first majority class distinctions, then the majority-minority split, and finally minority class differences.
-See Fig.
-13 for a visualization of $\bm{U}$ and Fig.
-9 for an experiment that confirms the above semantic interpretation of concepts (columns of $\bm{U}$).
-True ClassStep 0Refer to captionmaj 1maj 2min 1min 2maj 1maj 2min 1min 2Step 13Refer to captionmaj 1maj 2min 1min 2Step 30Refer to captionmaj 1maj 2min 1min 2Step 55Refer to captionmaj 1maj 2min 1min 2Step 60 Refer to captionmaj 1maj 2min 1min 2Predicted Class
-Figure 9: Experimental illustration of the fact that important concepts are learned first.
-Confusion matrix evolution during training of a 3-layer convolutional network on an imbalanced MNIST data.
-The matrices show five snapshots at different training steps (0, 13, 30, 55, 65) for four classes: two majority classes (Maj1, Maj2) with 100 samples each and two minority classes (Min1, Min2) with 10 samples each.
-Training progresses from left to right: initially classifying all data as Maj1 (Step 0), then correctly identifying majority classes while misclassifying minority classes (Step 13), gradually improving minority class recognition (Step 30), confusion between minority classes only (Step 55), and finally achieving perfect classification by Step 60.
-Appendix B From Sparsity Language Pattern to Concepts Geometry
-Recall the centered data-sparsity matrix $\widetilde{\bm{S}}$ and its SVD
+Finally, while drawing conceptual similarities to [Saxe et al., 2019], our work differs in its motivation, interpretation, and scope of results.
+Their focus is on general cognitive development using an abstract two-layer linear neural network with orthogonal inputs.
+In contrast, we specifically analyze how semantic and grammatical concepts emerge from natural language data through NTP optimization.
+Methodologically, we realize a connection and apply their framework to the NTP-UFM (which we interpret as a linear two-layer network where the input dimension equals the number of input examples), a tractable model for expressive large architectures.
+This provides a practical instantiation for their otherwise restrictive orthogonal-input assumptions.
+We also corroborate our findings with experiments on real datasets and architectures, a crucial distinction from their purely theoretical work.
+Appendix C Derivations and Proofs
+C.1 Concepts in Embedding Space
+This section includes derivations of the formulas of word-concept and context-concept representations of Sec. 3.1. Recall the centered data-sparsity matrix $\widetilde{\bm{S}}$ and its SVD
 $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top},\quad\text{where}\quad\bm{U}\in\mathbb{R}^{V\times r},\bm{\Sigma}\in\mathbb{R}^{r\times r},{\bm{V}}\in\mathbb{R}^{m\times r}\quad\text{and}\quad\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}\,$,
-and the singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered:
+with singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ that are ordered:
 $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}>0\,.$
-Adopting terminology from Saxe et al.
-[2019], we interpret the columns $\bm{u}_{k}\in\mathbb{R}^{V},{\bm{v}}_{k}\in\mathbb{R}^{m},k\in[r]$ of $\bm{U},{\bm{V}}$ as word and context analyzer vectors for concept $k$.
+We interpret the columns $\bm{u}_{k}\in\mathbb{R}^{V},{\bm{v}}_{k}\in\mathbb{R}^{m},k\in[r]$ of $\bm{U},{\bm{V}}$ as word and context analyzer vectors for concept $k$.
 For each word $z\in\mathcal{V}$ and each word-concept $k\in[r]$, the component $\bm{u}_{k}[z]$ represents how present or absent is a word $z$ in context $k$.
 Respectively for contexts.
-Specifically, we think of column dimensions of $\bm{U}$ as semantic dimensions that capture semantic categories.
+Thus, we think of column dimensions of $\bm{U}$ as concepts that capture semantic categories.
+For completeness, we note that each word-analyzer vector $\bm{u}_{k}$ has both positive and negative entries, as it is orthogonal to $\mathds{1}_{V}$ (which lies in the nullspace of $\widetilde{\bm{S}}$).
 Q: How do we define word-concept and context-concept representations, i.e. $d$-dimensional representations of word and context analyzer vectors for various concepts?
@@
 let
-$\bm{u}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{W}}{\bm{W}}^{\top}\bm{u}_{k}\,$
-${\bm{v}}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{H}}{\bm{H}}{\bm{v}}_{k}\,.$
-Let’s now simplify these by using the known SVD representation of ${\bm{W}}$ and ${\bm{H}}$.
+$\bm{u}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{W}}{\bm{W}}^{\top}\bm{u}_{k}\,\qquad\text{and}\qquad{\bm{v}}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{H}}{\bm{H}}{\bm{v}}_{k}\,.$
+We can further simplify these by using the known SVD representation of $\bm{W}$} and $\bm{H}$} from Sec.
+2.
 Using this representation (i.e.
-use ${\bm{W}}\leftarrow{\bm{W}}^{\rm{mm}}$, ${\bm{H}}\leftarrow{\bm{H}}^{\rm{mm}}$) we compute
-$\mathds{P}_{\bm{W}}=\mathbf{R}\mathbf{R}^{\top}=\mathds{P}_{\bm{H}}.$
+use ${\bm{W}}\leftarrow{\bm{W}}^{\rm{mm}}$, ${\bm{H}}\leftarrow{\bm{H}}^{\rm{mm}}$) we compute $\mathds{P}_{\bm{W}}=\mathbf{R}\mathbf{R}^{\top}=\mathds{P}_{\bm{H}},$ where recall that $\mathbf{R}$ is a partial orthogonal matrix.
 Thus,
@@
 The second concept (${\bm{v}}^{\rm{d}}_{2}$, green line) represents a continuum from non-mammalian to mammalian traits, with mammalian words and features (“cat”, “stays warm-blooded”) having positive projections and non-mammalian ones (“trout”, “has feathers”) having negative projections.
-Please note that the concepts are not orthogonal in the plots because it is PCA projection to 2D space.
-Appendix C On the hierarchical structure of language
-As discussed in Sec.
-4.3, language semantics exhibit a hierarchical structure, with more comprehensive semantics involving numerous concepts being inherently more detailed.
-Each concept’s significance is indicated by its associated singular values; higher values denote concepts critical for differentiating words or contexts.
-Moreover, as discussed in Sec.
-4.4, broader semantic categories are acquired more rapidly, indicating a sequential learning progression from general to more specific concepts.
-Fig.
-11 below provides an illustration of this hierarchical structure.
-Figure 11: Hierarchical semantic structure revealed by orthant-Based clustering on Simplified TinyStories.
-Top: Hierarchical semantic structure emerges as we combine concepts.
-Each row represents clusters formed by combining an increasing number of concept dimensions (from 1 to 5), with each word cloud displaying the most typical words for a specific sign configuration.
-Semantic categories become progressively more fine-grained moving downward as additional concept dimensions refine the classifications.
-Bottom: Detailed view of verb semantics.
-The combination of the first two concept dimensions ($\mathcal{K}=\{1,2\}$ with configuration $C_{\mathcal{K}}=\{+1,+1\}$) identifies verbs.
-Adding the third dimension further distinguishes between past tense ($C_{\mathcal{K}}=\{+1,+1,+1\}$) and present tense ($C_{\mathcal{K}}=\{+1,+1,-1\}$).
-The present tense category is further refined by the fifth dimension to differentiate present action verbs from present continuous (-ing) forms.
-This hierarchical organization mirrors linguistic grammar structures and supports our hypothesis that language models learn fundamental grammatical distinctions (verb vs.
-non-verb) before more specific ones (tense and aspect distinctions).
-Appendix D Connections to closed-form training dynamics Saxe et al. [2013]
+(The concepts are not orthogonal in the plots because it is PCA projection to 2D space.)
+C.2 Proof of Theorem 1
 Consider the following square-loss NTP-UFM proxy:
@@
 Most works focus on global minima of regularized UFM, with less attention to unregularized cases or training dynamics.
-While some landscape analyses provide partial answers about global convergence (e.g., ), they are limited to regularized cases and don’t characterize dynamics.
+While some landscape analyses provide partial answers about global convergence, they are limited to regularized cases and do not characterize dynamics.
 Even for square loss, where Mixon et al.
@@
 The key insight is rewriting (6) as $\sum_{j\in[m]}|\widetilde{\bm{s}}_{j}-{\bm{W}}{\bm{H}}\bm{e}_{j}|^{2}$ with orthogonal inputs $\bm{e}_{j}\in\mathbb{R}^{m}$.
-This enables direct application of their result, originally stated in Saxe et al.
+This enables application of their result, originally stated in Saxe et al.
 [2013] and formalized in Gidel et al.
 [2019].
-For completeness, we state this here in our setting and terminology as a corollary below.
-Proposition 1.
-Consider gradient flow (GF) dynamics for minimizing the square-loss NTP-UFM (5). Recall the SVD $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$. Assume weight initialization
+The following is a detailed (re)statement of Theorem 1.
+Theorem 2.
+Consider gradient flow (GF) dynamics for minimizing the square-loss NTP-UFM (6).
+Recall the SVD $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$.
+Assume weight initialization (see Fig.
+9 for empirical justification of this form of spectral initialization)
 ${\bm{W}}(0)=e^{-\delta}\bm{U}\mathbf{R}^{\top}\quad\text{and}\quad{\bm{H}}(0)=e^{-\delta}\mathbf{R}{\bm{V}}^{\top}$
@@
 ${\bm{W}}(t)\rightarrow{\bm{W}}_{\infty}:=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}^{\top}\qquad\text{and}\qquad{\bm{H}}(t)\rightarrow{\bm{H}}_{\infty}:=\mathbf{R}\sqrt{\bm{\Sigma}}{\bm{V}}^{\top},.$ (10)
-This aligns with Zhao et al. [2024]’s regularization-path analysis of NTP-UFM with CE loss, where normalized quantities converge as ${\lambda}\rightarrow 0$ (recall Section 3.2):
+This aligns with the regularization-path analysis of NTP-UFM with CE loss, discussed in Sec. 2, where normalized quantities converge as ${\lambda}\rightarrow 0$:
 $\overline{{\bm{W}}}({\lambda})\rightarrow\bm{U}\sqrt{\overline{\bm{\Sigma}}}\mathbf{R}^{\top}\qquad\text{and}\qquad\overline{{\bm{H}}}({\lambda})\rightarrow\mathbf{R}\sqrt{\overline{\bm{\Sigma}}}{\bm{V}}^{\top}.$
@@
 Middle-Left: Convergence of word and context gram-matrices and of logits to the theory predicted by Thm.
-1.
+2.
 Middle-Right: Convergence of singular values of logit matrix to those of $\bm{\Sigma}$ (see Thm.
-1.
+2).
 Right: Projection of logits to subspace orthogonal to $\bm{U}$ and $\bm{V}$}; Logits with Rand initialization initially have non-zero projection but it becomes zero as training progresses.
-D.1 An example of controlling the rate of learning via reweighting
-Consider minimizing the following weighted version of (5):
-$\min_{{\bm{W}},{\bm{H}}}~{}~{}\left\|\left(\widetilde{\bm{S}}-{\bm{W}}{\bm{H}}\right)\bm{\Omega}\right\|^{2}\,,$ (10)
-where $\bm{\Omega}=\operatorname{diag}([\omega_{1},\ldots,\omega_{m}])$ is a diagonal matrix of weights, one for each context.
-Here, we consider the STEP-imbalanced one-hot classification setting described in Sec.
-A.4.
-Concretely, let the support matrix be
-$\bm{S}=\begin{bmatrix}\mathds{I}_{V/2}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}_{V/2\times V/2}\\ \mathbf{0}_{V/2\times RV/2}&\mathds{I}_{V/2}\end{bmatrix}$ (11)
-where $R$ is the imbalance ratio and without loss of generality we assumed that the first $V/2$ classes are majorities and that minorities have $1$ example (in our language: context) each.
-(Thus, the total number of examples is $m=RV/2+V/2=(R+1)V/2$.) Recall that $\widetilde{\bm{S}}=(\mathds{I}_{V}-\frac{1}{V}\mathds{1}_{V}\mathds{1}_{V}^{\top})\bm{S}$.
-Recall from Sec. A.4 that $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ where the singular values and the left singular vectors are explicitly given by:
-$\bm{\Sigma}=\operatorname{diag}(\begin{bmatrix}\sqrt{R}\mathds{1}_{V/2-1}&\sqrt{(R+1)/2}&\mathds{1}_{V/2-1}\end{bmatrix}),\qquad\bm{U}=\left[\begin{array}[]{ccc}\mathbb{F}&-\sqrt{\frac{1}{V}}\mathds{1}&\mathbf{0}\\ \mathbf{0}&\sqrt{\frac{1}{V}}\mathds{1}&\mathbb{F}\end{array}\right]\in\mathbb{R}^{V\times(V-1)}\,,$
-with $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$. According to [Thrampoulidis et al., 2022, Lem. A.3] we also have
-${\bm{V}}^{\top}=\left[\begin{array}[]{ccc}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]\in\mathbb{R}^{(V-1)\times m}\,.$
-We now set the weight matrix as follows:
-[…]
-Direct calculation yields
-${\bm{V}}^{\top}\bm{\Omega}$ $=\left[\begin{array}[]{ccc}\sqrt{\frac{1}{R}}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{2}{R(R+1)V}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]$ (16)
-$=\operatorname{diag}(\begin{bmatrix}\sqrt{\frac{1}{R}}\mathds{1}_{V/2-1}^{\top}&\sqrt{\frac{2}{R+1}}&\mathds{1}_{V/2-1}^{\top}\end{bmatrix})\underbrace{\left[\begin{array}[]{ccc}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{1}{RV}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{1}{V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]}_{=:\widetilde{\bm{V}}^{\top}}=\bm{\Sigma}^{-1}\widetilde{\bm{V}}^{\top}\,.$ (20)
-where in the last equation, we recognized the diagonal matrix is equal to $\bm{\Sigma}^{-1}$ and called the other matrix $\widetilde{\bm{V}}\in\mathbb{R}^{m\times(V-1)}$.
-Thus, the effective sparsity matrix $\widetilde{\bm{S}}_{\bm{\Omega}}=:\widetilde{\bm{S}}\bm{\Omega}$ is equal to
-$\widetilde{\bm{S}}_{\bm{\Omega}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}\bm{\Omega}=\bm{U}\widetilde{\bm{V}}^{\top}\,.$
-It is now easy to check that $\widetilde{\bm{V}}^{\top}\widetilde{\bm{V}}=\mathds{I}_{V-1}$.
-Thus, the display above is the SVD factorization of $\widetilde{\bm{S}}_{\bm{\Omega}}$.
-Note that the singular values of this effective sparsity matrix are all equal, unlike the singular values $\bm{\Sigma}$ of the original data-sparsity matrix $\widetilde{\bm{S}}$.
-With this at hand, we can show analogous to Proposition 1 that the eigenvalues of ${\bm{W}}(t)$ and ${\bm{H}}(t)$ are now all learned at the same time $T=1$.
-We also confirm this experimentally in Fig.
-13.
-Figure 13: (Top) Support matrix and SVD factors of centered support matrix for a one-hot STEP-imbalanced example.
-(Middle) Training dynamics of GD minimization of NTP-UFM with square loss (Eq.
-(5); same setup and visualizations as Fig.
-12.
-(Bottom) Training dynamics of GD minimization of NTP-UFM with weighted square loss (Eq.
-(10)) with weights as in Eq.
-(12).
-Note that thanks to the weighting all singular factors are now learned simultaneuously.
-Appendix E Extended Discussion on Related Works
-Word embeddings and semantic analysis in neural probabilistic language models.
-The word2vec architecture Mikolov et al.
-[2013d, b] and its variants, notably GloVe Pennington et al.
-[2014], represent seminal early neural probabilistic language models.
-These simple log-bilinear models, trained on large text corpora, revolutionized word embedding learning.
-As noted in Zhao et al.
-[2024], NTP-UFM shares structural similarities with these early models, though in both their work and ours, it serves as a tractable abstraction rather than a practical architecture.
-Our approach differs by learning both context and word embeddings, following modern practice.
-The foundational work of Levy and Goldberg [2014] connected word2vec’s geometry to matrix factorization of the pointwise mutual information (PMI) matrix—a specialized word co-occurrence matrix.
-Subsequent works Levy et al.
-[2015], Turney and Pantel [2010], Baroni and Lenci [2010] empirically demonstrated semantic interpretations of the PMI matrix’s singular factors and principal components.
-Building on Zhao et al.
-[2024], which formalizes a modern version of Levy and Goldberg [2014]’s results, our investigation of concepts differs from this classical literature in two key aspects: (a) We study the NTP setting where both context and word embeddings are learned, yielding concepts that relate to both words and contexts; (b) Our data-sparsity matrix differs fundamentally from classical PMI matrices: it is a centered version of the data support matrix (independent of specific next-token probabilities) and has different structural properties—being orthogonal and non-square, unlike the word2vec setting.
-Superposition and feature steering.
-Our work was partly motivated by recent compelling literature suggesting that embeddings can be decomposed into linear combinations of a finite set of semantic concepts Bricken et al.
-[2023], Yun et al.
-[2023], Park et al.
-[2023].
-These insights from mechanistic interpretability have led to practical applications in "feature steering"—where model behavior can be controlled by manipulating concept representations through addition or subtraction Durmus et al.
-[2024], Konen et al.
-[2024].
-Our analysis complements the mechanistic interpretability approach by providing a systematic framework for understanding how concepts emerge naturally as principal components from training data statistics.
-Exploring deeper connections between our theoretical framework and the mechanistic interpretability literature remains an intriguing direction for future work.
-For completeness, we note that Park et al.
-[2023, 2024] also investigate geometric properties of concept directions, albeit through fundamentally different technical approaches, assumptions, and perspectives.
-Saxe et al.’s closed-form dynamics of two-layer linear network training.
-Our work draws inspiration from Saxe et al.
-[2019].
-Conceptually, Saxe et al.
-[2019] uses a two-layer linear neural network as a theoretical proxy to study the emergence of semantic knowledge in human cognition, providing mathematical justifications for phenomena observed in cognitive semantics literature.
-A key insight from their work is that even a simple two-layer linear network with orthogonal inputs can yield rich and meaningful conclusions about semantic learning.
-While two-layer neural networks represent perhaps the simplest instances of non-linear learning, their training dynamics generally remain analytically intractable.
-However, Saxe et al.
-[2013] (with aspects later formalized in Gidel et al.
-[2019]) demonstrated that with square loss, orthogonal inputs, and sufficiently small initialization, these dynamics admit exact closed-form solutions.
-This mathematical characterization underlies their results on semantic information development through singular factors of the network’s input-output correlation matrix.
-We make a novel connection to this line of work: the UFM fits perfectly within the framework studied by Saxe et al.
-[2013, 2019].
-Specifically, the UFM can be viewed as a linear two-layer network where the input dimension equals the number of input examples (in our case, the number of contexts $m$).
-This connection is valuable in two directions: First, the UFM—recently popularized through neural collapse literature (see below)—provides perhaps the most natural and practical setting satisfying Saxe et al.’s seemingly restrictive orthogonal input assumptions.
-Second, this connection allows us to leverage Saxe et al.’s earlier results in the evolving neural collapse literature.
-Despite these methodological similarities with Saxe et al.
-[2019], our work differs in motivation and interpretation.
-We focus specifically on NTP and how semantic and grammatical concepts emerge from natural language data, rather than general cognitive development.
-Additionally, we primarily focus on CE loss which is typically used in NTP training.
-Neural-collapse geometries.
-Our results contribute to the recent literature on the neural collapse (NC) phenomenon Papyan et al.
-[2020a].
-Originally observed in one-hot classification training of DNNs, neural collapse describes two key properties of well-trained, sufficiently expressive DNNs: (1) NC: embeddings of examples from the same class collapse to their class mean, and (2) ETF-geometry: class-mean embeddings form a simplex equiangular tight frame (ETF), being equinorm and maximally separated, with classifier weights exhibiting the same structure and aligning with their respective class-mean embeddings.
-This phenomenon, consistently observed across diverse datasets and architectures, has sparked extensive research interest, generating hundreds of publications a complete review of which is beyond our scope.
-Instead, we discuss below the most closely-related line of works.
-One fundamental direction, which forms the basis for many extensions, focuses on explaining NC’s emergence through the UFM.
-This model abstracts training as joint optimization of last-layer embeddings (unconstrained by architecture) and classifier weights Mixon et al.
-[2022], Fang et al.
-[2021].
-Multiple influential works have proven NC emergence by analyzing the UFM’s global optima (e.g., Zhu et al.
-[2021], Garrod and Keating [2024], Jiang et al.
-[2023]), with extensions to various loss functions beyond cross-entropy, including square loss Zhou et al.
-[2022a], Han et al.
-[2021], Tirer and Bruna [2022], Súkeník et al.
-[2023, 2024] and supervised contrastive loss Graf et al.
-[2021], Zhou et al.
-[2022b], Kini et al.
-.
-Most early works maintained the original assumptions from Papyan et al.
-[2020a]: balanced data (equal examples per class) and embedding dimension $d$ exceeding the number of classes $C$.
-Recent work has explored $d<C$ settings, though often requiring additional assumptions on the loss function Jiang et al.
-[2023], Liu et al.
-[2023].
-More substantial progress has emerged in the $d>C$ regime with unbalanced data, where Thrampoulidis et al.
-[2022] provided a complete characterization for step-imbalanced data (where examples are distributed equally within minority classes and equally within majority classes).
-They introduced the SELI (simplex-encoding labels interpolation) geometry, showing that logits interpolate a simplex-encoding matrix—a centered version of the one-hot encoding matrix.
-The embeddings and classifier vectors are then determined, up to rotation and scaling, by the singular vectors of this matrix.
-The SELI geometry emerges as a special case of the richer geometries characterized in the NTP setting by Zhao et al.
-[2024].
-Together with Li et al.
-[2023], these works stand alone in extending geometric characterization beyond one-hot encoding—to soft-label and multilabel settings respectively.
-Specifically, Zhao et al.
-[2024] analyzes the soft-label setting arising in NTP training on natural language, showing that word and context embeddings are determined by the singular factors of the data sparsity matrix.
-Our work deepens this understanding by revealing that these SVD factors encode conceptual meaning, thereby extending neural collapse geometry to capture not only the structure of embeddings but also the organization of latent concepts.
-Appendix F Additional Experiment Results
-F.1 Semantics for Individual Concept
-In this section we include the plots illustrating that individual concept do not contain interpretable semantics as discussed in sec 5
+Appendix D Additional Experimental Results
+D.1 Minimal One-Hot Imbalanced Setting
+In Sec. 3.3, we identified class imbalance as the minimal requirement for progressive semantic emergence. Here, we provide detailed analysis of this minimal setting.
+The emergence of semantically meaningful concepts stems from the structure of labels encoded in $\bm{S}$ (or its centered version $\widetilde{\bm{S}}$) across different contexts.
+Consider the minimal case: each context has exactly one next-token, with contexts distributed equally across the vocabulary.
+Here, $\bm{S}$ can be rearranged as $\mathds{I}_{V}\otimes\mathds{1}_{m/V}^{\top}$, yielding trivial concepts where all singular directions contribute equally.
+This setting corresponds to balanced one-hot classification studied in the neural collapse literature, where last-layer embeddings and weights form highly symmetric aligned structures Papyan et al.
+[2020a], reflecting the symmetric nature of $\bm{S}$ with no interesting conceptual structure.
+However, such balanced settings are unrealistic for language modeling.
+As seen in the controlled and real experiments in Secs.
+3 and 4, natural language produces rich label formations where $\bm{S}$ induces a geometry of concepts that provides semantic interpretation to embedding structures.
+Here, we demonstrate that meaningful concepts emerge even in one-hot classification—the traditional focus of neural collapse literature—given minimal deviation from perfect balance.
+Specifically, consider class imbalance with ratio $R$: $V/2$ majority classes each have $R>1$ times more samples than the $V/2$ minority classes.
+It can be analytically shown that $\widetilde{\bm{S}}$ has singular values exhibiting a three-tier structure [Thrampoulidis et al., 2022b]:
+$\sigma_{1}$ $=\ldots=\sigma_{V/2-1}=\sqrt{R}$ (11)
+$>\sigma_{V/2}=\sqrt{(R+1)/2}$ (12)
+$>\sigma_{V/2+1}=\ldots=\sigma_{V-1}=1.$ (13)
+Moreover, the left singular vectors matrix $\bm{U}$ takes a sparse block form:
+$\bm{U}=\left[\begin{array}[]{ccc}\mathbb{F}&-\sqrt{\frac{1}{V}}\mathbf{1}&\mathbf{0}\\ \mathbf{0}&\sqrt{\frac{1}{V}}\mathbf{1}&\mathbb{F}\end{array}\right]\in\mathbb{R}^{V\times(V-1)},$
+where $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ is an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$.
+[^4] For concreteness, $\mathbb{F}$ can be constructed using the discrete cosine transform matrix, excluding the constant column: $\mathbb{F}[i,j]=\sqrt{\frac{4}{V}}\cdot\cos\left(\frac{\pi(2i-1)j}{V}\right)$ for $i\in[V/2]$, $j\in[V/2-1]$.
+The structure of $\bm{U}$ reveals three distinct concept types, corresponding to the three tiers of singular values:
+(1) First V/21V/2-1 columns (largest singular values): Non-zero only for majority classes, representing distinctions among majority classes.
+(2) Middle column (singular value $\sqrt{(R+1)/2}$): Opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy.
+(3) Last V/21V/2-1 columns (smallest singular values): Non-zero only for minority classes, capturing distinctions among minority classes.
+This structure demonstrates that the network learns concepts in order of singular value magnitude: first majority class distinctions, then the majority-minority split, and finally minority class differences.
+Fig.
+10 shows an experiment on imbalanced MNIST that confirms this semantic interpretation of concepts (columns of $\bm{U}$) and validates that class imbalance alone suffices for progressive semantic emergence.
+True ClassStep 0Refer to captionmaj 1maj 2min 1min 2maj 1maj 2min 1min 2Step 13Refer to captionmaj 1maj 2min 1min 2Step 30Refer to captionmaj 1maj 2min 1min 2Step 55Refer to captionmaj 1maj 2min 1min 2Step 60 Refer to captionmaj 1maj 2min 1min 2Predicted Class
+Figure 10: Experimental illustration of the fact that important concepts are learned first.
+Confusion matrix evolution during training of a 3-layer convolutional network on an imbalanced MNIST data with $V=4$ classes, 2 majorities, 2 minorities and imbalance ratio $R=10$.
+The matrices show five snapshots at different training steps (0, 13, 30, 55, 65) for four classes: two majority classes (Maj1, Maj2) with 100 samples each and two minority classes (Min1, Min2) with 10 samples each.
+Training progresses from left to right: initially classifying all data as Maj1 (Step 0), then correctly identifying majority classes while misclassifying minority classes (Step 13), gradually improving minority class recognition (Step 30), confusion between minority classes only (Step 55), and finally achieving perfect classification by Step 60.
+For this setting, Thrampoulidis et al.
+[2022a] show the singular values of $\widetilde{\bm{S}}$ exhibit a three-tier structure: $\sigma_{1}=\ldots=\sigma_{V/2-1}=\sqrt{R}>\sigma_{V/2}=\sqrt{(R+1)/2}>\sigma_{V/2+1}=\ldots=\sigma_{V}=1.$ Additionally, The structure of $\bm{U}$ reveals three distinct types of concepts, corresponding to the three tiers of singular values: (1) First V/21V/2-1 columns (non-zero only for majority classes) represent distinctions among majority classes; (2) Middle column (singular value $\sqrt{(R+1)/2}$) has opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy; (3) Last V/21V/2-1 columns (non-zero only for minority classes) capture distinctions among minority classes.
+Thus, the evolution of confusion matrices in the figure perfectly aligns with this hierarchical structure.
+D.2 Training transformer on Simplified Tinystories dataset
+For Simplified TinyStories, we train 6-layer and a 12-layer decoder-only transformers, with 8 attention heads, and 1024 hidden dimensions.
+Training uses the AdamW optimizer with learning rate 1e-3, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay 0.001.
+We use batch size 64 and train until convergence (approximately 1000 epochs).
+D.3 On the hierarchical structure of language
+As discussed in Sec.
+3.2, language semantics exhibit a hierarchical structure, with more comprehensive semantics involving numerous concepts being inherently more detailed.
+Fig.
+11 below provides an illustration of this hierarchical structure.
+Figure 11: Hierarchical semantic structure revealed by orthant-based clustering on Simplified TinyStories.
+Top: Clusters are formed by combining 1 to 5 concept dimensions; each row shows word clouds for different sign configurations.
+Finer-grained semantic categories emerge as more dimensions are added.
+Bottom: Verb-related structure: $\mathcal{K}={1,2}$ highlights verbs; adding a third dimension separates past and present tense; the fifth dimension further distinguishes verb aspects.
+D.4 Semantics for Individual Concept
+We include the plots illustrating that individual concept do not contain interpretable semantics as discussed in Sec 4.
 Figure 12: Word clouds of top words from individual concept dimensions across datasets, illustrating their lack of human-interpretable semantics without combination.
 Shown: positive 2nd dimension (Simplified TinyStories), negative 6th dimension (Simplified TinyStories), negative 2nd dimension (Simplified WikiText), positive 4th dimension (Simplified WikiText).
-F.2 Experiment on the Semantics of Pretrained Context Embeddings
-To identify the semantics in context embeddings from pretrained transformers, we chose GPT-2 Radford et al.
-[2019] and extracted $50,000$ text sequences, each 11 tokens in length from GPT-2’s training corpus, the OpenWebText dataset.
-By segmenting these sequences into contexts of varying lengths (1-10 tokens), we generated $500,000$ contexts, which we designate as the "simplified OpenWebText" dataset.
-To obtain concepts from pretrained models, instead of performing SVD on $\widetilde{\bm{S}}$ (which is computationally expensive to construct and factorize), we work directly with the embeddings of the pretrained model.
-For word concepts, we extract ${\bm{W}}$ from GPT-2’s decoder and take the matrix $\bm{U}$ of word concepts to be its left singular matrix.
-For context concepts, we form a $d\times m$ matrix ${\bm{H}}$ by concatenating the embeddings (last-MLP layer representations before the decoder) of GPT-2 for each context in our subset.
-This matrix corresponds to only a portion of the “true” embedding matrix for the entire dataset.
-To ensure the coupling between ${\bm{W}}$ and ${\bm{H}}$ that is inherent to the SVD of $\widetilde{\bm{S}}$ (which we don’t have direct access to), we compute the matrix ${\bm{V}}$ of context concepts as ${\bm{V}}={\bm{H}}^{\top}\mathbf{R}\bm{\Sigma}^{-1/2}$, where $\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ denotes the SVD of ${\bm{W}}$ that we already computed.
-We include the orthant-clustering result on GPT2’s pretrained word embedding with $p=4,5,6$. (Fig. 15, 16, 17)
-Figure 15: orthant-based clustering on Pretrained word embedding, $p=4$, GPT2
-Figure 16: Semantics on Pretrained word embedding, $p=5$, GPT2
-Figure 17: orthant-based clustering on Pretrained word embedding, $p=6$, GPT2
-F.3 Rate of learning
+D.5 Rate of learning
 Fig. 13 shows the evolution of singular values during training for both squared loss and cross-entropy loss as described in Section 3.3.
@@
 (Right) With CE loss, singular values grow unboundedly while maintaining their relative ordering, reflecting the continuous growth of embedding norms characteristic of CE training.
-F.4 $d>V$
-Fig. 19 shows the evolution of singular values for squared loss when the network’s embedding dimension is smaller than the vocabulary size, i.e., $d<V$, as described in Section A.1.
-Figure 19: evolution of the learned logit matrix’s singular values during training for $3=d<V$, note that the 3 singular values are converging to the $d$ largest singular values
-F.5 Semantics from K-means Spectral Clustering on Simplified TinyStories and Wiki-Text
-Please see Fig. 20 and 21.
-Figure 20: 32-means spectral clustering, $p=5$, Simplified TinyStories
-Figure 21: 64-means spectral clustering, $p=6$, Simplified WikiText
-F.6 More orthant-based clustering result on Simplified TinyStories and Wiki-Text
-We include some examples of orthant-based clustering results in Fig.
-22 to 25.
-We selected $p=4$, $p=5$ for Simplified TinyStories and $p=5$, $p=6$ for Simplified WikiText because we observe the most human-interpretable concepts.
-Figure 22: Orthant-based clustering, $p=4$, Simplified TinyStories
-Figure 23: Orthant-based clustering, $p=5$, Simplified TinyStories
-Figure 24: Orthant-based clustering, $p=5$, Simplified Wiki-Text
-Figure 25: Orthant-based clustering, $p=6$, Simplified Wiki-Text
+D.6 Extension to Masked Language Models
+To explore semantic structure in model trained with different objectives, we applied orthant-based clustering to token embeddings from a pretrained BERT model.
+As shown in Fig.
+14, the method recovers a range of interpretable semantic categories.
+3-digit Number
+Historical Years
+Action
+Administration
+Given Names
+Operations
+Name Suffixes
+Verb Suffixes
+Figure 14: Semantics identified by orthant-based clustering on BERT’s word embeddings. In BERT’s tokenizer, tokens prefixed with ## represent subword units that appear within or at the end of a word.
+D.7 Extension to Higher-Dimensional Embeddings
+As discussed in Sec. 4.2, we extend our analysis to the Qwen-7B model. We summarize representative semantic categories for selected orthant configurations in Table 1.
+Table 1: Examples of semantic categories recovered from Qwen-7B using orthant-based clustering on its 4096-dimensional word embeddings.
+Each row corresponds to a specific sign configuration $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ along the top singular concept directions.
+D.8 Semantics from K-means Spectral Clustering on Simplified TinyStories and Wiki-Text
+Unlike orthant-based clustering, which can produce up to $2^{p}$ clusters when selecting $p$ top concepts, vanilla k-means on principal components results in $p$ clusters.
+Thus, when evaluating spectral methods, we choose $k$ on the order of $2^{p}$ for fair comparison.
+Specifically We apply $k$-means clustering using pairwise distances of analyzer vectors restricted to their top $p=\log_{2}(k)$ dimensions.
+As mentioned, this choice corresponds to the top-$p$ concept directions anddiffers from vanilla spectral clustering, which selects $k$ principal components.
+We chose $k=32$ for Simplified TinyStories and $k=64$ for WikiText-2.
+Fig.
+15 displays example clusters.
+Past-tense verbs Refer to caption
+Proper names Refer to caption
+Numbers Refer to caption
+Conjunctions Refer to caption
+Figure 15: $k$-means cluster with top-$p=\log_{2}(k)$ dimensions of word-analyzer vectors.
+Left: 2 of 32-means clusters with $\mathcal{K}=\{1,\ldots,5\}$ on Simplified TinyStories representing Past-tense verbs and Proper names; Right: 2 of 64-means cluster with $\mathcal{K}=\{1,\ldots,6\}$ on simplified WikiText, representing Numbers and Conjunctions.
+Appendix E Use of AI Assistants
+We used AI assistants (ChatGPT, Claude, and GitHub Copilot) to support manuscript preparation.
+ChatGPT and Claude assisted with text editing, grammar correction, and phrasing refinement.
+GitHub Copilot assisted with code generation and debugging.
+All conceptual contributions, theoretical results, experimental designs, and scientific conclusions are solely the work of the authors.
+AI-generated content was critically reviewed by the authors.